### v1

# Information Elicitation in Agency Games

Abstract
Rapid progress in scalable, commoditized tools for data collection and data processing has made it possible for firms and policymakers to employ ever more complex metrics as guides for decision-making. These developments have highlighted a prevailing challenge—deciding which metrics to compute. In particular, a firm’s ability to compute a wider range of existing metrics does not address the problem of unknown unknowns, which reflects informational limitations on the part of the firm. To guide the choice of metrics in the face of this informational problem, we turn to the evaluated agents themselves, who may have more information than a principal about how to measure outcomes effectively. We model this interaction as a simple agency game, where we ask: When does an agent have an incentive to reveal the observability of a cost-correlated variable to the principal? There are two effects: better information reduces the agent’s information rents but also makes some projects go forward that otherwise would fail. We show that the agent prefers to reveal information that exposes a strong enough differentiation between high and low costs. Expanding the agent’s action space to include the ability to garble their information, we show that the agent often prefers to garble over full revelation. Still, giving the agent the ability to garble can lead to higher total welfare. Our model has analogies with price discrimination, and we leverage some of these synergies to analyze total welfare.
1 Introduction
The rise of algorithmic and data-driven decision-making has come with an accompanying rise in reliance on numerical measurements of performance. These numerical metrics drive actions at scales ranging from that of individual workers such as teachers [Koretz, 2017], to entire institutions such as hospitals [Muller, 2019], and can have far-reaching impact on social welfare. The problem of designing and choosing evaluation metrics thus continues to be both highly consequential and challenging for all types of organizations, from government agencies to companies.
For example, consider the evaluation-metric design problem faced by a company with a research division. To monitor the health of the organization and guide compensation, metrics are collected to evaluate the research division’s productivity. To choose these metrics, company leadership may start with metrics obtained from ad hoc brainstorming or prior experience, taking the form of simple measures such as the number of academic publications, patents, conference presentations, and citations to all of these, along with collaborations with product teams. Such measures may not, however, fully capture the performance of the research division—in fact, they will almost inevitably omit some important factors. The challenge in designing a monitoring system is thus not only in developing methods for measuring performance and optimizing a given set of metrics, but more broadly how to bring new metrics to the fore that were previously unknown and are highly relevant.
This frames our central question of how to improve a firm’s specification of metrics. The perspective underlying this work is that while company leadership may operate under incomplete information, this does not mean that better information does not exist elsewhere in the firm. In fact, often, agents being evaluated may have better information not only about their effort, but also about how to measure outcomes in a more effective, less gameable way. In the above example, researchers might know of other metrics capturing the quality of their work that company leadership had overlooked, like the comparative quality of publication venues. Thus, we model and analyze situations under which an agent being evaluated might be willing to share information with the principal to improve the metrics that are used for evaluation. A key strategic dependence in this work is the relationship between the informant (the agent) and the metrics—we consider an agent’s incentives to reveal information about metrics that are directly used to evaluate them.
The incomplete nature of metrics has been discussed in the seminal work of Holmström and Milgrom [1991] on contracts with multidimensional tasks, where a principal may observe only a subset of dimensions that are relevant to their value or the agent’s cost. When the principal is aware of which dimensions are missing, Holmström and Milgrom [1991] show how to optimally reward the observed dimensions given properties of the agent’s cost structures. The distinguishing feature of our setting is that we consider the unobserved dimensions to be unknown unknowns. We focus on an information transfer mechanism where the agent has the power to possibly reveal these hidden dimensions to the principal.
Our question also fits into a broad class of problems on information design and persuasion games that studies the welfare effects of different information structures [Kamenica and Gentzkow, 2011, Milgrom and Roberts, 1986]. Our model treats an agent as a sender and principal as a receiver, and can be seen as applying specific restrictions to the class of information-revelation strategies that the agent can choose from in order to capture properties of the design of metrics in agency problems faced by firms. Specifically, we consider a setting where the agent’s information revelation strategy cannot depend on the value of the realized signal. Thus, we seek to model an agent’s choice of whether or not to reveal the observability of a metric, rather than a realized signal value.
To yield initial tractable insights to this metric discovery problem, we examine an agent’s incentives for information sharing through the lens of an agency game with information transfer. We build on a classical agency game where a principal contracts an agent to complete a task, and the principal only has partial information about the agent’s costs when setting a contract. To capture the agent’s additional information and opportunity to improve the metrics by which they are evaluated, our model supposes that the agent is privately aware of additional variables that correlate with their cost of task completion, and further has the opportunity to reveal these additional variables to the principal prior to the design of the contract. We analyze when the agent would prefer for the contract to depend on these cost-correlated variables compared to an agnostic contract. Importantly, the agent must decide whether to reveal their cost-correlated variables to the principal prior to realizing the values of those variables or their true costs.
While the principal is always better off when having more information about the agent’s costs, the incentives for the agent are more nuanced—revealing information reduces the amount of information rent the agent can extract when their costs turn out to be low. However, better information also means more high-cost jobs will go forward that the agent otherwise would not have accepted, as the optimal contract would adequately reward high-cost tasks.
We first consider whether the agent prefers to reveal or conceal the observability of a cost-correlated environmental variable. We show that the agent prefers to reveal the variable if conditioning on this cost-correlated variable reveals a strong enough differentiation between high and low costs of task completion. Next, we expand the agent’s action space to include the ability to garble or add noise to their information before it passes to the principal. For example, in modern online platforms, avenues for introducing this type of noise include third party clients or policies requiring differential privacy Dwork et al. [2006]. We show that under a fairly wide set of conditions, the agent may prefer to reveal a garbled signal over both fully concealing and fully revealing the original variable. This suggests that a noisy information transfer mechanism can yield superior equilibria over simpler or more restrictive alternatives.
1.1 Contributions
Our contributions can be summarized as follows:
  1. 1.
    We introduce a model for the process of discovery of unknown metrics, in the form of an agency game with information transfer.
  2. 2.
    We present sufficient conditions under which an agent would prefer revealing a metric to a principal over keeping the metric concealed; and vice versa.
  3. 3.
    We further relax the agent’s action space to include the ability to reveal garbled information, where the choice of the amount of garbling interpolates between concealment and revelation. In this setting, we give sufficient conditions for the agent to prefer garbling over full revelation.
  4. 4.
    We analyze the consequences of information revelation on principal utility and total welfare, leveraging connections between our model and price discrimination.
1.2 Related Work
Our model builds on literature from contract design and agency games. It also fits into a large literature on information design, and can be seen as a specific structure of information design problem. Our model also overlaps with price discrimination, and we bring techniques from these literatures to a new motivation of discovering metrics.
Agency games and contract design
We build from the well-established contract design problem of Laffont and Tirole [1986], which concerns a principal’s design of a contract when an agent’s effort and cost type are privately held by the agent. Key to this setting are asymmetries in values and information between the principal and agent, and the literature explores issues of moral hazard and adverse selection that arise from these asymmetries Laffont and Martimort [2009], Gibbons et al. [2013], Milgrom and Roberts [1992].
A fundamental result regarding signaling incentives in contract theory is Holmström [1979]’s sufficient statistic theorem, which showed that it benefits an agent for the contract to be conditioned on any information that is independently informative of the agent’s effort. Our setting models an agent’s incentive to share information about its cost type, which yields different results from the analysis of effort signaling, and is closer to some analyses of persuasion games which we discuss in more detail below. Milgrom [1981] also classifies the “favorableness” of signals to an agent, presenting monotonicity properties that we also leverage in this work.
Persuasion games and information design
The use of stakeholder-supplied information for decision-making was seminally introduced by Milgrom and Roberts [1986] in the form of persuasion games. Milgrom and Roberts [1986] give an example of understanding a buyer’s purchasing strategy in the face of a seller who can send a quality signal about their product. While this broad motivation of information transfer from interested parties is very close to our motivation of learning business metrics from evaluated agents, the game that we present differs in two main ways from the particular setting in Milgrom and Roberts [1986].
First, we restrict the agent’s information revelation strategy for a given variable to not be able to depend on the realized value of the variable. In other words, the agent must commit to fully revealing a variable in advance of realizing the variable’s value. In this sense, our setting can be viewed as a constrained persuasion game. The purpose of this restriction is to bring our setting closer to a problem of variable discovery, where the fundamental problem is a principal’s lack of awareness of a variable’s observability, rather than signaling Spence [1978], where the uncertainty is about the variable’s realized value.
The second main distinction is in the direction of information transfer. In our setting, the “decision-maker” is the agent, who decides whether or not to complete a task. We consider the agent’s incentive to reveal information about their decision-making parameters (their cost type) to a principal who designs a payment contract. This second distinction is more technical and less important for the fundamental motivation of revealing observability instead of value.
More broadly, Bayesian persuasion and information design provides a general framework for analyzing the effects of the distribution of information on the outcomes of a game Kamenica [2019], Bergemann and Morris [2019]. Many works have cast this general form to analyze applications from grading systems in schools to courtroom evidence policies Kamenica and Gentzkow [2011], Boleslavsky and Cotton [2015], Ostrovsky and Schwarz [2010]. Our model can be seen as a specific instantiation of an information design problem where the sender is the agent, the receiver is the principal, and the contracting relationship determines the principal’s and agent’s action spaces and equilibria. We impose a constraint over the sender’s information transfer policy in order to capture the incentives for an agent to reveal observability of a variable to the principal, whereas prior work has focused on revealing the value of a signal (similarly to Milgrom and Roberts [1986]). Also built into this constraint is that the agent cannot lie about their signal, as our model is motivated by understanding settings where a principal has powerful data collection and verification capabilities.
Price discrimination
Our analysis framework of an agency game with information transfer has analogies with classical price discrimination, and thereby opens new avenues for applying the well-developed tools from price discrimination to understanding the metric design problem. It also uncovers new challenges that extend existing price discrimination perspectives. This connection has also been discussed by Bergemann et al. [2015], who define a mapping from third degree price discrimination onto the class of agency problems and establish the existence of market segmentations that achieve all possible trade-offs between consumer and producer surplus within some basic constraints. Our work complements this analysis—instead of analyzing all possible segmentations, we consider the agent’s information-revelation incentives as a function of the properties of a specific segmentation induced by some cost-correlated variable which is initially only available to the agent.
Our garbling setting also notably differs from the models of price discrimination with transportation costs and arbitrage [Wright, 1993] or restricted price discrimination Aguirre et al. [2010]. While both garbling and restricted price discrimination mechanisms effectively interpolate between full discrimination and no discrimination, we prove a substantive difference between these mechanisms by showing that while the agent might prefer some intermediate amount of garbling over both full revelation or hiding, the agent will always prefer either full discrimination or no discrimination over any intermediate restriction over the amount of allowed price discrimination. We give a more detailed comparison of our garbling model to price discrimination in Section 4.7.
Sunspots and correlated equilibria
Our model is also connected to the so-called sunspots literature Woodford [1990], Howitt and McAfee [1992]. In this literature, there are multiple steady-state equilibria and an observable random variable produces a correlated equilibrium, with agents conditioning their behavior on this otherwise extraneous variable not because it matters to payoffs, but because it predicts others’ behaviors. The literature is known as sunspots because, during the 19th century, some people believed sunspot activity predicted agricultural yields [Jevons, 1884], and while it did not, it could predict the behavior of commodity traders who believed it did. This literature is the polar opposite of the problem we study: sunspots are a known, extraneous variable believed to be relevant, while we study an unknown (to the principal) relevant variable.
Preferences for privacy
Our model also interacts with literatures on privacy and information-hiding, by highlighting a simple situation where agents experience conflicting incentives both for and against sharing information. The result can be that agents may prefer partial sharing, which echoes the subtleties that arise elsewhere when privacy and behavior interact Cummings et al. [2016].
2 Model Setup: Agency Game with Information Transfer
To gain insight into the incentives surrounding the discovery of metrics, we start with a standard agency game in which a principal contracts an agent to complete a task. In such a game, we ask, when does the agent have an incentive to reveal observability of a cost-correlated variable to the principal?
Specifically, suppose a principal contracts the agent to complete a task, where an agent may exert binary effort. Suppose the principal receives value $b$ if the agent exerts effort and completes the task, and zero otherwise (we assume that the task is completed deterministically if the agent exerts effort). Suppose the agent incurs cost $C\in\mathbb{R}_{+}$ for exerting effort. The exact cost is unobserved to the principal, but the principal is aware of a prior distribution over agent’s cost type, denoted by the random variable $C$. The agent observes both the proposed transfer and their realized cost type before deciding whether or not to exert effort. This maps onto the well-established agency game setup where the principal must design a contract when the agent’s cost and effort are private [Laffont and Tirole, 1986].
To model a setting where the agent might possess additional information, suppose the agent is aware of an environmental variable $X\in\mathcal{X}$ which is correlated with cost $C$. The question of knowledge of $X$ becomes significant when $X$ is correlated with $C$.
In introducing this correlation into the agency game, we assume that $X$ is an unknown unknown to the principal. Prior to the principal’s design of the contract, the agent has a choice of whether or not to inform the principal of the environmental variable $X$, which entails revealing both the observability of $X$ at the time of the design of the contract, and the realized value of $X$ at the execution of the contract (which the principal can verify). If the agent chooses not to inform the principal of the observability of $X$, then the rest of the game proceeds as a standard agency game with private cost: the principal designs a contract based on their knowledge of the prior distribution $C$. If the agent chooses to inform the principal of $X$, then the principal can offer a contract that conditions on the realized value of $X$.
The timing of the full agency game with the possibility of agent information transfer is summarized in Figure 1. The text in black matches the standard agency game [Laffont and Tirole, 1986], and the text in gray represents additional elements introduced by our model.
$t=0$ P and A share prior distribution over $C$. A knows prior joint distribution of $C,X$. $t=1$ A decides whether or not to reveal observability of $X$, including the joint distribution of $C,X$. $t=2$ P offers a contract, which can depend on $X$ if observability of $X$ was revealed. $t=3$ $X$ is realized. A learns their cost type $C$. $t=4$ A decides whether or not to accept the contract. $t=5$ The contract is executed and utilities realized.
Figure 1: Timing of the agency game with information transfer between principal (P) and agent (A).
In our simplified environment involving binary effort, the principal’s contract design problem when $X$ is concealed reduces to choosing a single transfer amount $p$ where the agent is paid $p$ if the task is completed, and zero otherwise. If $X$ is revealed, then the principal offers a contract with distinct transfers $\rho(x)$ for different realized values of $x\in\mathcal{X}$. At the time of the execution of the contract, the agent receives transfer $\rho(x)$ if the task is completed and $X=x$; and zero otherwise. We assume that the principal still receives the same value $b$ if the task is completed, regardless of $X$.
The key question in this work concerns the agent’s decision of whether or not to inform the principal of the existence of the environmental variable $X$ at time $t=1$. In Section 3, we consider this as a binary decision of whether or not to reveal $X$; we will later relax this in Section 4 to expand the agent’s action space to reveal a garbled version of $X$, thus interpolating between the concealed and revealed settings.
2.1 Notation
Let $F(c)={\mathbb{P}}(C\leq c)$ denote the cumulative distribution function (CDF) of the prior cost distribution $C$. We also define $F_{x}(c)={\mathbb{P}}(C\leq c|X=x)$ as the CDF of the conditional distribution of $C$ given $X$. Throughout, we will assume that the cost $C$ is a continuous random variable with density $f(c)=F^{\prime}(c)$ and with conditional density $f_{x}(c)=F_{x}^{\prime}(c)$.
We will use $\Pi$ to denote the principal’s utility, $V$ to denote the agent’s utility, and $W$ to denote total welfare, defined as the sum of the principal’s and agent’s utilities.
Let $\operatorname{\mathbbm{1}}(\cdot)$ denote an indicator function with $\operatorname{\mathbbm{1}}(x\in S)=\begin{cases}1&\text{if }x\in S\\ 0&\text{otherwise}.\end{cases}$
All proofs are given in the Appendix.
3 Welfare Effects of Information Revelation
We now consider the agent’s decision of whether to conceal or reveal the existence $X$ at time $t=1$ in the agency game with information transfer outlined in Figure 1. To analyze this decision, we compare the agent’s utility in the information-agnostic contract that results from keeping $X$ concealed and in the informed contract that depends on $X$. We also analyze the consequences of the resulting decision on the principal’s utility and on total welfare.
As an overview, we give sufficient conditions for the agent to prefer to conceal, and sufficient conditions for the agent to prefer to reveal $X$. In general, the principal always prefers the revealed setting. Finally, we analyze the consequences of revelation on total welfare, connecting to analogous results from price discrimination.
3.1 Utilities for Principal and Agent
We begin by outlining the contract and equilibrium behavior of the principal and agent when the contract is agnostic of $X$, which we refer to as the concealed information setting. Then, we outline the contract and equilibrium behavior when the principal can condition on $X$ to determine payments to the agent, which we refer to as the revealed information setting. We assume that both principal and agent are risk neutral throughout.
3.1.1 Concealed information contract
If information is not revealed, then the rest of the game proceeds as a standard agency game with private cost, where the principal chooses a single transfer $p$ based on their knowledge of the prior distribution over the agent’s cost $C$. The agent’s optimal policy at the execution of the contract is to exert effort if their realized cost is less than or equal to the payment. Thus, the agent’s best response to the principal’s choice of transfer $p$ is to exert effort with probability $F(p)={\mathbb{P}}(C\leq p)$ (the set where $C=p$ has measure zero).
For a given choice of transfer $p$, the principal’s expected utility when $X$ is hidden is given by
$\Pi_{\text{con}}(p)\coloneqq F(p)(b-p).$ (1)
The utility of the agent under the principal’s choice of transfer $p$ is given by
$V_{\text{con}}(p)\coloneqq\mathbb{E}[(p-C)\operatorname{\mathbbm{1}}(C<p)].$ (2)
The principal moves first and chooses $p^{*}\in\arg\max_{p\geq 0}\Pi_{\text{con}}(p)$. The agent’s utility at equilibrium is then $V_{\text{con}}(p^{*})$.
3.1.2 Revealed information contract
If the environmental variable $X$ is revealed at time $t=1$, then the principal has the ability to instead choose a transfer which depends the value of $X$, denoted as $\rho\mathrel{\mathop{\ordinarycolon}}\mathcal{X}\to\mathbb{R}_{+}$. Since the agent is aware of the values of both $C$ and $X$ before deciding whether to exert effort, and there is no possibility to lie about $X$, the agent’s optimal policy is to exert effort if their realized cost is less than or equal to the transfer given the realized $X$ value. Therefore, the agent’s best response to the principal’s transfer function $\rho(\cdot)$ is to exert effort with probability $F_{x}(\rho(x))$ when $X=x$. The principal’s expected utility in this revealed setting is given by
$\Pi_{\text{rev}}(\rho)\coloneqq\mathbb{E}[F_{X}(\rho(X))(b-\rho(X))]=\mathbb{E}[\Pi_{X}(\rho(X))],$
where $\Pi_{x}(p)\coloneqq F_{x}(p)(b-p)$.
The utility of the agent under the principal’s choice of transfer function $\rho(\cdot)$ is then
$V_{\text{rev}}(\rho)\coloneqq\mathbb{E}[(\rho(X)-C)\operatorname{\mathbbm{1}}(C<\rho(X))]=\mathbb{E}[V_{X}(\rho(X))],$
where $V_{x}(p)\coloneqq\mathbb{E}[(p-c)\operatorname{\mathbbm{1}}(C<p)|X=x]$.
For ease of exposition, our analysis will focus on a one-dimensional binary feature: $\mathcal{X}=\{0,1\}$, where $X=1$ with probability $\theta$. We frame the principal’s decision problem as that of choosing $\rho(0)=p_{0}$ and $\rho(1)=p_{1}$. Further extensions to higher cardinalities or continuous $X$ may be of practical interest, though we expect our qualitative insights to extend to these settings.
The principal’s expected utility becomes
$\Pi_{\text{rev}}(p_{0},p_{1})\coloneqq(1-\theta)\Pi_{0}(p_{0})+\theta\Pi_{1}(p_{1}),$
and the agent’s utility becomes
$V_{\text{rev}}(p_{0},p_{1})\coloneqq(1-\theta)V_{0}(p_{0})+\theta V_{1}(p_{1}).$
The principal moves first and chooses $p_{0}^{*},p_{1}^{*}\in\arg\max_{p_{0},p_{1}}\Pi_{\text{rev}}(p_{0},p_{1})$, resulting in an equilibrium where the agent’s utility is $V_{\text{rev}}(p_{0}^{*},p_{1}^{*})$.
Without loss of generality, we will refer to the situation when $X=1$ as the “stronger” situation with generally lower cost for effort. That is, $p_{0}^{*}>p_{1}^{*}$.
In the rest of this section, we will analyze the effects of information revelation on the agent’s utility, principal’s utility, and total welfare. Specifically, we will compare $V_{\text{con}}(p^{*})$ to $V_{\text{rev}}(p_{0}^{*},p_{1}^{*})$ to understand the agent’s incentive to reveal $X$. We will also analyze the consequences of the agent’s revelation decision on the principal’s utility, as well as consequences on total welfare.
3.2 Agent’s Revelation Incentives
The central goal in this work is to analyze the circumstances under which the agent would prefer to either conceal or reveal the existence of $X$ at time $t=1$. Thus, to begin, we will analyze situations that lead to $V_{\text{con}}(p^{*})$ being higher than $V_{\text{rev}}(p_{0}^{*},p_{1}^{*})$ or vice versa, focusing on properties of the distributions $F$, $F_{0}$, and $F_{1}$. In the analogy to price discrimination, we may think of $F(p)$ as the task completion “quantity” as a function of price $p$, or the proportion of agents drawn uniformly at random from a population with costs distributed as $C$ that would complete the task for price $p$.
We present sufficient conditions on the distributions $F_{0},F_{1}$ for the agent to either prefer to conceal or to prefer to reveal. We then build intuition for these conditions through an example using exponential and Weibull distributions.
3.2.1 Concealment condition with one zero-cost type
First, we present a sufficient condition for the agent to prefer to conceal $X$ when one of the agent types is anchored at zero. That is, $X=1$ implies that the agent incurs zero cost. Proposition 1 gives a sufficient condition on $F_{0}$ for the agent to prefer for the environmental variable $X$ to remain concealed.
Proposition 1 (Sufficient concealment condition with zero-cost type).
Suppose $F_{0}$ is a concave and continuously differentiable CDF. Suppose $C|X=1$ takes value 00 with probability $1$. Suppose the ratio $\frac{F_{0}(p)}{f_{0}(p)}$ is strictly monotone increasing for $p>0$. Then $V_{\text{con}}(p^{*})>V_{\text{rev}}(p_{0}^{*},p_{1}^{*})$ if
$\theta>(1-\theta)\frac{1}{\eta((1-\theta)p_{0}^{*})}-\frac{1}{\eta_{0}\left(p_{0}^{*}\right)},$ (3)
where $\eta(p)=\frac{p(1-\theta)f_{0}(p)}{(1-\theta)F_{0}(p)+\theta}$ and $\eta_{0}(p)=\frac{pf_{0}(p)}{F_{0}(p)}$ are the respective price elasticities for task completion quantity for the mixture distribution $C$ and the conditional distribution $F_{0}$.
Qualitatively, the elasticity $\eta$ captures the sensitivity of the task completion to price. Thus, the inequality in equation (3) corresponds to a scenario when the sensitivity of the task completion to price when $X=0$ does not differ too strongly from that of the concealed setting. For example, this arises when $F_{0}$ is close to the constant function 1. We give another example using the exponential distribution in Section 3.3 below, where equation (3) holds if the mean of $F_{0}$ is low enough. In summary, the agent prefers concealment if the higher cost type still has relatively low cost.
3.2.2 Concealment and revelation conditions under a decreasing ratio assumption
To give additional sufficient conditions for concealment and revelation beyond the anchored setting with one zero-cost type, we apply an analysis technique similar to that of Aguirre et al. [2010], who analyzed the effects of third degree monopoly price discrimination on total welfare.
Suppose the principal, on knowing $X$, is constrained to choose transfers $p_{0},p_{1}$ subject to the constraint that $p_{0}-p_{1}<r$ for some $r\geq 0$. Let $p_{0}(r),p_{1}(r)$ denote the principal’s optimal transfers under this constraint:
$\begin{split}p_{0}(r),p_{1}(r)\in&\arg\max_{p_{0},p_{1}}\;\;\Pi_{\text{rev}}(p_{0},p_{1})\\ &\text{s.t.}\;\;p_{0}-p_{1}\geq r.\end{split}$ (4)
For notational convenience, let $V_{\text{const}}(r)\coloneqq V_{\text{rev}}(p_{0}(r),p_{1}(r))$. In a similar structure to Aguirre et al. [2010], the results in this section come from considering the “marginal effect of relaxing the constraint” on the agent’s value.
Under the following closely analogous assumptions to those invoked by Aguirre et al. [2010], we derive properties of $V_{\text{const}}(r)$.
Assumption 1 (Concave principal utility).
The principal’s utility in each realized environment is strictly concave: $\Pi_{0}^{\prime\prime}(p)<0$, $\Pi_{1}^{\prime\prime}(p)<0$.
Assumption 2 (Decreasing ratio condition (DRC)).
The ratios $\frac{V_{0}^{\prime}(p)}{\Pi_{0}^{\prime\prime}(p)}$ and $\frac{V_{1}^{\prime}(p)}{\Pi_{1}^{\prime\prime}(p)}$ are both decreasing in $p$.
Assumption 2 is analogous to the “increasing ratio condition” assumption from Aguirre et al. [2010], which instead has the derivative of total welfare in the numerator. Our analysis naturally extends this to focus on agent utility. Assumption 2 holds in almost the same set of conditions as the assumption on total welfare from Aguirre et al. [2010], and we discuss the subtleties of the differences between these assumptions in Appendix A.3.
Lemma 1.
Under Assumptions 1 and 2, $V_{\text{const}}(r)$ is strictly quasi-convex for $r\in[0,p_{0}^{*}-p_{1}^{*}]$. That is, if there exists $\hat{r}\in[0,p_{0}^{*}-p_{1}^{*}]$ such that $V_{\text{const}}(\hat{r})=0$, then $V_{\text{const}}^{\prime\prime}(\hat{r})>0$.
The strict quasi-convexity of the agent’s utility in $r$ makes it possible to derive sufficient conditions for revelation and concealment by differentiating $V_{\text{const}}$ and evaluating the sign of the derivative at extreme values of $r$. Adapting this machinery from Aguirre et al. [2010], but focusing on the agent’s value instead of total welfare, we give such sufficient conditions for the agent to prefer concealing or revealing $X$ below.
Proposition 2 (Sufficient concealment condition under DRC).
Under Assumptions 1 and 2, $V_{\text{con}}(p^{*})>V_{\text{rev}}(p_{0}^{*},p_{1}^{*})$ if
$\frac{(1-\theta)(b-p_{0}^{*})}{2-\sigma_{0}(p_{0}^{*})}<\frac{\theta(b-p_{1}^{*})}{2-\sigma_{1}(p_{1}^{*})},$ (5)
where $\sigma_{x}(p)=\frac{F_{x}(p)f^{\prime}_{x}(p)}{f_{x}^{2}(p)}$ is the curvature of the inverse of the task completion quantity function $F_{x}(p)$.
Proposition 2 implies that a high enough difference in curvature between $\sigma_{0}(p_{0}^{*})$ and $\sigma_{1}(p_{1}^{*})$ implies that the agent will prefer the concealed contract over the revealed contract. That is, inverse task completion quantity when $X=0$ is more convex than the inverse task completion quantity when $X=1$ at the revealed transfers $p_{0}^{*},p_{1}^{*}$. This is exactly the flipped version of the condition in Aguirre et al. [2010]’s Proposition 2, which implied that total welfare is higher under price discrimination. We next give a sufficient condition for the agent to prefer revelation.
Proposition 3 (Sufficient revelation condition under DRC).
Under Assumptions 1 and 2, $V_{\text{con}}(p^{*})<V_{\text{rev}}(p_{0}^{*},p_{1}^{*})$ if
$\frac{2+L(p^{*})\alpha_{1}(p^{*})}{2+L(p^{*})\alpha_{0}(p^{*})}>\frac{\theta F_{1}(p^{*})/f_{1}(p^{*})}{(1-\theta)F_{0}(p^{*})/f_{0}(p^{*})},$ (6)
where $L(p)=\frac{b-p}{p}$ is the Lerner index Lerner [1995], and $\alpha_{x}(p)=\frac{-pf_{x}^{\prime}(p)}{f_{x}(p)}$ is the curvature of the task completion quantity function $F_{x}(p)$,
Intuitively, Proposition 3 says that if the curvatures of $F_{0}$ and $F_{1}$ are different enough (relative to the ratio of the CDFs themselves), then the agent will prefer to reveal the environmental variable $X$.
Remark
An important property analyzed by Milgrom [1981] is the monotone likelihood ratio property (MLRP), which here would say that $\frac{f_{0}(c)}{f_{1}(c)}$ is increasing for all $c$. The MLRP would imply that both sides of the inequality in equation (6) are greater than 1. However, this does not necessarily imply an order between these ratios, and there exist distributions that satisfy the MLRP that yield either of the inequality directions above. We give a specific example of this using the Weibull distribution in Section 3.3 below.
3.3 Example: Exponential and Weibull Distributions
To concretely illustrate the conditions in Propositions 1, 2, and 3, we parameterize the conditional cost distributions using the exponential distribution and the more general Weibull distribution.
First, to illustrate the condition in Proposition 1, let $C|X=0\sim\text{Exp}(\frac{1}{\lambda_{0}})$, where $\lambda_{0}$ represents the scale parameter and is also the mean of the distribution. Specifically,
$F_{0}(c)=\begin{cases}1-e^{-\frac{1}{\lambda_{0}}c}&c\geq 0\\ 0&c<0.\end{cases}$ (7)
Then the condition in Equation (3) is equivalent to $\lambda_{0}<b\psi(\theta)$, where $\psi(\theta)=\frac{\theta}{\left(\left(\frac{1}{1-\theta}\right)^{\frac{1}{\theta}}-\left(\frac{1}{1-\theta}\right)^{\frac{1-\theta}{\theta}}-\theta\right)}$ is monotone decreasing function bounded between 00 and $1$ for $\theta\in[0,1]$. Thus, as long as the average cost $\lambda_{0}$ is less than a $\theta$-dependent scaling of the task completion value $b$, the condition in Proposition 1 holds. In other words, if the average cost when $X=0$ is not too high, then the agent will prefer concealment.
Beyond fixing $F_{1}$ at zero cost, let $C$ be a mixture of exponential distributions with $C|X=0\sim\text{Exp}(\frac{1}{\lambda_{0}})$ and $C|X=1\sim\text{Exp}(\frac{1}{\lambda_{1}})$, where
$F_{x}(c)=\begin{cases}1-e^{-\frac{1}{\lambda_{x}}c}&c\geq 0\\ 0&c<0.\end{cases}$ (8)
Figure 2 plots the difference $V_{\text{rev}}(p_{0}^{*},p_{1}^{*})-V_{\text{con}}(p^{*})$ for all $\lambda_{0},\lambda_{1}\in[0,b]$. As seen in Proposition 1, if $\lambda_{1}=0$, then the agent prefers to hide if $\lambda_{0}$ is sufficiently low. This continues to hold for $\lambda_{1}$ sufficiently close to 0. More generally, Figure 2 shows that the agent prefers to reveal if the means $\lambda_{0},\lambda_{1}$ are sufficiently far apart.
$V_{\text{rev}}(p_{0}^{*},p_{1}^{*})-V_{\text{con}}(p^{*})$
$\lambda_{1}$
$\lambda_{0}$
Figure 2: Difference between agent’s utility in the revealed setting and concealed settings when $C$ is distributed as a mixture of exponentials. For each pair $(\lambda_{0},\lambda_{1})$, a positive value indicates that the agent prefers revelation, and a negative value indicates that the agent prefers concealment. The contour line shows all $(\lambda_{0},\lambda_{1})$ for which $V_{\text{rev}}(p_{0}^{*},p_{1}^{*})-V_{\text{con}}(p^{*})=0$. The parameters $b=1$ and $\theta=\frac{1}{2}$ are fixed, and $\lambda_{0}$ and $\lambda_{1}$ are varied up to $b$.
For the exponential mixture, the inequalities in Propositions 2 and 3 do not hold for any combinations of $\lambda_{0},\lambda_{1}$. Thus, the condition in Proposition 1 covers cases not covered by Proposition 2. However, we see Propositions 2 and 3 take effect for the more general Weibull distribution, with
$F_{x}(c)=\begin{cases}1-e^{\left(-\frac{1}{\lambda_{x}}c\right)^{k_{x}}}&c\geq 0\\ 0&c<0.\end{cases}$ (9)
For example, for a fixed $\lambda_{0},\lambda_{1}$, increasing $k=k_{0}=k_{1}$ increases the difference in curvature between $F_{0}$ and $F_{1}$ at $p^{*}$, yielding a set of values of $k>1$ in which the condition in Proposition 3 holds. For Proposition 2, the condition holds if $k_{1}$ is sufficiently small, and $k_{0}$ is sufficiently large for $\lambda_{0}>\lambda_{1}$.
3.4 Principal’s Revelation Preferences
While the agent might sometimes prefer the hidden setting over the revealed setting, we next show that the principal always prefers revelation. First, Lemma 2 shows that the principal is never worse off under revelation.
Lemma 2 (Principal prefers revelation).
Revealing $X$ never decreases the value of the principal: $\Pi_{\text{rev}}(\rho^{*})\geq\Pi_{\text{con}}(p^{*})$, where $\rho^{*}\in\arg\max_{\rho}\Pi_{\text{rev}}(\rho)$. Revealing $X$ strictly increases the value of the principal only if $X$ and $C$ are not independent.
The principal strictly benefits from information revelation if the monotone likelihood ratio property (MLRP) is satisfied between the revealed distributions.
Assumption 3 (Monotone likelihood ratio property (MLRP) [Milgrom, 1981]).
The ratio $\frac{f_{0}(c)}{f_{1}(c)}$ is strictly increasing in $c$.
Lemma 3 (Principal strictly benefits from revelation).
Let $F_{0}$ and $F_{1}$ be continuously differentiable CDFs. If the MLRP holds (Assumption 3), then the principal strictly benefits when $X$ is revealed: $\Pi_{\text{rev}}(p_{0}^{*},p_{1}^{*})>\Pi_{\text{con}}(p^{*})$.
Intuitively, as the first mover, the principal will never be hurt by having additional freedom to condition on $X$ when selecting prices that maximize their utility. Lemma 3 gives the MLRP as a sufficient condition for revelation of $X$ to yield a strict benefit for the principal.
3.5 Total Welfare Consequences
We now consider whether revealing $X$ increases total welfare, or the sum of utilities of the principal and agent. When $X$ is concealed, total welfare is given by
$W_{\text{con}}(p)\coloneqq V_{\text{con}}(p)+\Pi_{\text{con}}(p),$
and when $X$ is revealed, total welfare is given by
$W_{\text{rev}}(p_{0},p_{1})\coloneqq V_{\text{rev}}(p_{0},p_{1})+\Pi_{\text{rev}}(p_{0},p_{1}).$
The question of whether price discrimination increases total welfare has been well studied Varian [1985]. Our comparison of the concealed vs. revealed contracts has a direct isomorphism with third degree monopoly price discrimination. Specifically, our concealed setting corresponds to monopoly pricing without price discrimination, with the seller acting as principal and buyer acting as agent. Our revealed setting corresponds to a monopoly seller enacting third degree price discrimination over markets segmented by $X$.
Thus, with minor adjustments, we can apply results from the price discrimination literature that characterize the effects of third degree monopoly price discrimination on total welfare. Mirroring Varian [1985]’s seminal work, Lemma 4 shows that total welfare increases only if the quantity of tasks completed also increases in the revealed setting compared to the concealed setting.
Lemma 4.
Total welfare increases under revelation ($W_{\text{rev}}(p_{0}^{*},p_{1}^{*})-W_{\text{con}}(p^{*})$) only if task completion quantity increases under revelation,
$F(p^{*})<(1-\theta)F_{0}(p_{0}^{*})+\theta F_{1}(p_{1}^{*}).$
3.5.1 Example when total welfare decreases
It is still possible for total welfare to decrease under information revelation. Mirroring an example from Varian [1985], we provide an illustrative example here where total welfare decreases when task completition quantity does not increase.
Suppose $C|X=1\sim\text{Unif}(0,1)$, and $C|X=0\sim\text{Unif}(\frac{1}{2},\frac{3}{2})$. Suppose $\theta=\frac{1}{2}$. Suppose $b=1$. Then the optimal payments for each of $F,F_{0},F_{1}$ all fall in the “interior” of $F(x)$:
$\frac{1}{2}\leq p_{1}^{*}<p^{*}<p_{0}^{*}\leq 1.$
When the solutions all fall in the interior, we have $p^{*}=\frac{p_{1}^{*}+p_{0}^{*}}{2}$, and $F(p^{*})=\frac{F_{1}(p_{1}^{*})+F_{0}(p_{0}^{*})}{2}$. This now violates the necessary condition in Lemma 4, since the output does not increase under revelation, but the payments change. Total welfare decreases as long as $p_{0}^{*}\neq p_{1}^{*}$.
4 Information Revelation With Garbling
So far, we have compared the settings when the environmental variable $X$ is either concealed or and revealed, focusing on the agent’s incentive to induce each setting at time $t=1$. We now generalize the agent’s action space to instead be able to reveal a garbled version of the variable $X$. From a modeling standpoint, the agent’s garbling action space interpolates between the concealed and revealed settings. In this work, we consider a randomized response garbling mechanism which has been applied in many settings, from statistical informativeness Blackwell [1951] to survey experiment design Warner [1965] to differential privacy Dwork et al. [2006], Kasiviswanathan et al. [2011].
As an overview of results in this section, we show that there exist conditions under which the agent would prefer to garble over both concealment and revelation. Thus, having the option to garble can benefit the agent, and even induce the agent to reveal more information when they would otherwise fully opt to conceal $X$. We also analyze the effects of garbling on the principal’s utility and total welfare.
4.1 Model Setup: Information Transfer with Garbling
Suppose the agent has the option to present the principal with a different variable $Y$ with random noise on top of $X$, defined as
$Y=\begin{cases}X&\text{w.p. }\varepsilon\\ \xi&\text{w.p. }1-\varepsilon\end{cases}=\begin{cases}X&\text{w.p. }\frac{1+\varepsilon}{2}\\ \neg X&\text{w.p. }\frac{1-\varepsilon}{2},\end{cases}$ (10)
where $\xi\sim\text{Bernoulli}(\gamma)$ is independent of $X$ and $C$. This work considers this randomized response mechanism for binary $X$, though further extensions with different noise models can be made for continuous $X$.
The game with garbling proceeds as before, but the agent selects $\varepsilon\in[0,1]$ at time $t=1$. The full timing is outlined in Figure 3. The agency game with garbled information transfer is a generalization of the previous game: in the previous timing in Figure 1, the agent’s choice at $t=1$ would be equivalent to selecting $\varepsilon$ from a more restricted set $\{0,1\}$.
$t=0$ P and A share prior distribution over $C$. A knows prior joint distribution of $C,X$. $t=1$ A selects $\varepsilon$, and reveals the observability of $Y$ (10), including joint distribution of $C,Y$. $t=2$ P offers a contract, which depends on $Y$. $t=3$ $X$ and $Y$ are realized. A learns their cost type $C$. $t=4$ A decides whether or not to accept the contract. $t=5$ The contract is executed and utilities realized.
Figure 3: Timing of the agency game with garbled information transfer between principal (P) and agent (A).
It is assumed that the principal, on knowing $Y$, cannot back out the previous value of $X$. The principal treats the variable $Y$ as an environmental variable with prior joint distribution $Y,C$, and proceeds to design the optimal contract conditioning on $Y$. The principal does not have access to $\varepsilon$ when designing the contract—only the variable $Y$ and its potential values.
The game proceeds as delineated for the revealed information contract in Section 3.1.2, but using the variable $Y\in\mathcal{Y}$ instead of $X$. Specifically, the principal designs a contract with transfer function $\rho\mathrel{\mathop{\ordinarycolon}}\mathcal{Y}\to\mathbb{R}_{+}$, and upon execution of the contract, the agent receives transfer $\rho(y)$ if they exert effort and $Y=y$. For binary $X$ and $Y$ following equation (10), the principal chooses $\rho(0)=p_{0}$ and $\rho(1)=p_{1}$.
Let the conditional distribution of the cost $C$ given $Y$ be distributed with CDF ${\mathbb{P}}(C\leq c|Y=y)=G_{y}(c)$. We focus our analysis on a simple randomized response noise model that does not change the marginal distribution over $X$, setting $\gamma=\theta=\frac{1}{2}$. In this case, we have
$G_{0}(c)=\frac{1+\varepsilon}{2}F_{0}(c)+\frac{1-\varepsilon}{2}F_{1}(c),\quad\text{and}\quad G_{1}(c)=\frac{1+\varepsilon}{2}F_{1}(c)+\frac{1-\varepsilon}{2}F_{0}(c).$
Similar results can be derived for general $\gamma,\theta$, and there is further room to explore the interaction between these parameters in future work.
As in Section 3.1, the utility of the principal under revelation of a given garbled variable $Y$ defined by equation (10) is given by
$\Pi_{\text{garb}}(p_{0},p_{1})=(1-\theta)E[(b-p_{0})\operatorname{\mathbbm{1}}(C<p_{0})|Y=0]+\theta E[(b-p_{1})\operatorname{\mathbbm{1}}(C<p_{1})|Y=1],$
and the utility of the agent is given by
$V_{\text{garb}}(p_{0},p_{1})=(1-\theta)E[(p_{0}-C)\operatorname{\mathbbm{1}}(C<p_{0})|Y=0]+\theta E[(p_{1}-C)\operatorname{\mathbbm{1}}(C<p_{1})|Y=1].$
The principal moves first and chooses $p_{0}(\varepsilon),p_{1}(\varepsilon)\in\operatorname*{arg\,max}_{p_{0},p_{1}}\Pi_{\text{garb}}(p_{0},p_{1})$. The agent’s resulting utility is $V_{\text{garb}}(p_{0}(\varepsilon),p_{1}(\varepsilon))$.
To simplify notation, we make the following abbreviations in the rest of this section: $\Pi_{\text{garb}}(\varepsilon)=\Pi_{\text{garb}}(p_{0}(\varepsilon),p_{1}(\varepsilon))$, $V_{\text{garb}}(\varepsilon)=V_{\text{garb}}(p_{0}(\varepsilon),p_{1}(\varepsilon))$.
4.2 Garbling and Prices
First, we analyze the effects of the information revelation amount $\varepsilon$ on the equilibrium prices set by the principal. Specifically, we show that both $p_{0}(\varepsilon),p_{1}(\varepsilon)$ are monotone with respect to $\varepsilon$.
Lemma 5 (Monotonic price changes).
Suppose $\gamma=\theta=\frac{1}{2}$. Suppose the principal’s utility is strictly concave as a function of price (Assumption 1). Suppose the MLRP holds (Assumption 3). Then $p_{0}^{\prime}(\varepsilon)>0$ and $p_{1}^{\prime}(\varepsilon)<0$ for all $\varepsilon\in[0,1]$.
Lemma 5 acts as a sanity check that as the principal is aware of more information, the degree of differentiation between prices also increases. Furthermore, with less noise, the price in the higher-cost environment will only increase, and the price in the lower-cost environment will only decrease.
4.3 Agent’s Garbling Incentives
For the agent, there is no clear ordering between $V_{\text{garb}}(p_{0}(\varepsilon),p_{1}(\varepsilon))$, $V_{\text{con}}(p^{*})$, and $V_{\text{rev}}(p_{0}^{*},p_{1}^{*})$. Garbling can be better for the agent than concealment, even when hiding is preferred over revelation, $V_{\text{con}}(p^{*})>V_{\text{rev}}(p_{0}^{*},p_{1}^{*})$. Garbling can also be better than revelation, even when revelation is preferred over hiding, $V_{\text{con}}(p^{*})<V_{\text{rev}}(p_{0}^{*},p_{1}^{*})$. We illustrate both of these cases using exponential distributions in Section 4.4, as previously done for the full concealment and revelation comparison in Section 3.3.
To give a more general theoretical characterization of the agent’s garbling incentives, we give sufficient conditions for the agent to prefer a nonzero amount of garbling over full revelation. This can be seen as a softer version of the previous analysis of when the agent prefers full concealment over full revelation.
4.3.1 Garbling condition under one zero-cost type
As in Section 3.2.1, we first analyze the restricted case where one of the agent types is anchored at zero-cost: suppose $C|X=1$ takes value 00 with probability one. Proposition 4 gives a sufficient condition for the agent to prefer a non-zero amount of garbling over full revelation.
Proposition 4 (Sufficient garbling condition with zero-cost type).
Suppose $\gamma=\theta=\frac{1}{2}$. Suppose $C|X=1$ takes value 0 with probability $1$. Suppose $F_{0}$ is continuously differentiable and $f_{0}(c)$ is bounded. $V_{\text{garb}}(\varepsilon)$ is maximized at $\varepsilon^{*}<1$ if
$\frac{b-p_{0}^{*}}{2-\sigma_{0}(p_{0}^{*})}<g_{0}(p_{0}^{*}),$ (11)
[…]
Notably, the inequality in equation (11) captures distributions that are not captured by Proposition 1. Thus, comparing Proposition 4 to Proposition 1 shows that the agent might want to garble, even if they may not always want to hide. In fact, the condition in equation (11) is quite general, and we show in the example in Section 4.4 below that equation (11) applies to any log-concave Weibull distribution.
4.3.2 General garbling condition
Generalizing beyond the anchored setting, Proposition 5 gives a sufficient condition for the agent to prefer garbling over revelation for general $F_{0},F_{1}$, which depends on similar identities. First, we generalize the restricted mean cost function $g_{0}(p)$ to a comparison of agent utilities.
Definition 1 (Agent utility dominance).
Let $\Delta(p_{0},p_{1})\coloneqq(V_{1}(p_{0})-V_{1}(p_{1}))-(V_{0}(p_{0})-V_{0}(p_{1}))$ denote the difference in sensitivities to the price change from $p_{0}$ to $p_{1}$ in each environment.
When $F_{1}$ exhibits first order stochastic dominance over $F_{0}$, we have that $\Delta(p_{0},p_{1})>0$ for $p_{0}>p_{1}$. The greater the dominance of $V_{1}(p)$ over $V_{0}(p)$ for all $p$, the greater the difference $\Delta$. Thus, we refer to $\Delta(p_{0},p_{1})$ as agent utility dominance. Using this definition, we now generalize the sufficient garbling condition from the anchored setting.
Proposition 5 (Sufficient garbling condition).
Suppose $\gamma=\theta=\frac{1}{2}$. Suppose $F_{0},F_{1}$ are continuously differentiable. $V_{\text{garb}}(\varepsilon)$ is maximized at $\varepsilon^{*}<1$ if
$-\Pi_{1}^{\prime}(p_{0}^{*})\left(\frac{b-p_{0}^{*}}{2-\sigma_{0}(p_{0}^{*})}\right)-\Pi_{0}^{\prime}(p_{1}^{*})\left(\frac{b-p_{1}^{*}}{2-\sigma_{1}(p_{1}^{*})}\right)<\Delta(p_{0}^{*},p_{1}^{*}),$ (12)
where $\sigma_{x}(p)=\frac{F_{x}(p)f_{x}^{\prime}(p)}{f_{x}(p)^{2}}$ is the curvature of the inverse quantity function.
The left hand side of the inequality in equation (12) is a weighted version of the difference $\frac{b-p_{0}^{*}}{2-\sigma_{0}(p_{0}^{*})}-\frac{b-p_{1}^{*}}{2-\sigma_{1}(p_{1}^{*})}$, which arises repeatedly in Aguirre et al. [2010]’s analysis of the effects of price discrimination on total welfare, and also previously arose in Proposition 2. In cases where $\Pi_{0}^{\prime}(p_{1}^{*})>-\Pi_{1}^{\prime}(p_{0}^{*})$, the condition in Proposition 2 (equation (3)) would imply the condition in Proposition 5 (equation (12)), since $\Delta(p_{0}^{*},p_{1}^{*})\geq 0$.
4.4 Example: Exponential and Weibull Distributions
We first illustrate the condition in Proposition 4 using an exponential distribution. Suppose $C|X=1$ takes value 00 with probability one, and suppose $C|X=0\sim\text{Exp}(\frac{1}{\lambda_{0}})$, with $F_{0}$ defined as in equation (7). In this case, $g_{0}(p_{0}^{*})=\lambda_{0}F_{0}(p_{0}^{*})$, and $\frac{b-p_{0}^{*}}{2-\sigma_{0}(p_{0}^{*})}=\lambda_{0}F_{0}(p_{0}^{*})\frac{1}{2-F_{0}(p_{0}^{*})}$. Therefore, the inequality in equation (11) holds for all $\lambda_{0}>0$.
The significance of this example is that if one agent type is anchored at 00, and the non-zero-cost environment induces an exponential distribution, the agent will always have an incentive to garble, regardless of the mean of the non-zero-cost distribution. Consider this in comparison to the exponential example from Section 3.3, where the condition in Proposition 1 showed that agent prefers to fully conceal $X$ when $\lambda_{0}$ is small enough.
For a Weibull distribution with $F_{0}$ given by equation (9), we have for $k_{0}\geq 1$,
$g_{0}(p)=\frac{\lambda}{k_{0}}\left(\Gamma\left(\frac{1}{k_{0}}\right)-\Gamma\left(\frac{1}{k_{0}},\frac{p^{k_{0}}}{\lambda_{0}^{k_{0}}}\right)\right).$
If $k_{0}\geq 1$, then the inequality in equation (11) from Proposition 4 holds for all $\lambda_{0}$. This encompasses all log-concave Weibull distributions. If $k_{0}<1$, then equation (11) does not necessarily hold, and fully flips for $k_{0}<0.5$.
Similarly to Figure 2, we can also consider simulations beyond the zero-cost anchored setting by considering all combinations of $\lambda_{0},\lambda_{1}$ when $C|X=0\sim\text{Exp}(\frac{1}{\lambda_{0}})$ and $C|X=1\sim\text{Exp}(\frac{1}{\lambda_{1}})$, with $F_{x}$ given by equation (8). Figure 4 illustrates the combinations of $\lambda_{0},\lambda_{1}$ for which the agent prefers some amount of garbling over full revelation. Figure 4 shows that $V_{\text{garb}}^{\prime}(1)<0$ as long one of the conditional means $\lambda_{0}$ or $\lambda_{1}$ is small enough. That is, as long as one of the revealed settings has low enough cost, the agent always prefers to garble, regardless of how high the cost of the other setting goes. This contrasts the revelation example in Section 3.3, where even if $\lambda_{1}$ is close to 00, high enough $\lambda_{0}$ leads to the agent being willing to reveal.
$V_{\text{garb}}^{\prime}(1)$
$\lambda_{1}$
$\lambda_{0}$
Figure 4: Plot of $V_{\text{garb}}^{\prime}(1)$ for a mixture of exponential distributions with means $\lambda_{0},\lambda_{1}$. A negative value indicates that the agent prefers some amount of garbling over full revelation. As long as one of the revealed settings has low enough average cost, the agent always prefers some amount of garbling over full revelation, regardless of how high the mean is of the other setting.
Finally, Figure 5 show a case in which the agent would prefer to reveal some amount of information via garbling over both concealment and revelation, but if not given the option to garble, then they would otherwise prefer concealment over revelation in the game from Figure 1.
4.5 Principal’s Garbling Preferences
Similarly to Section 3.4, the principal always prefers for more information to be revealed. In fact, the garbling parameter $\varepsilon$ directly interpolates between the concealed and revealed utilities for the principal.
Lemma 6.
$\Pi_{\text{con}}(p^{*})\leq\Pi_{\text{garb}}(p_{0}(\varepsilon),p_{1}(\varepsilon))\leq\Pi_{\text{rev}}(p_{0}^{*},p_{1}^{*})$ for all $\varepsilon$.
Also similarly to the comparison of full concealment and revelation settings, revealing less noisy information yields a strict improvement in the principal’s utility if the MLRP holds.
Lemma 7 (Strict principal improvement).
Suppose $\gamma=\theta=\frac{1}{2}$. Suppose the principal’s utility is strictly concave (Assumption 1). If the MLRP holds (Assumption 3), then $\Pi_{\text{garb}}^{\prime}(\varepsilon)>0$ for all $\varepsilon\in[0,1]$.
4.6 Garbling and Total Welfare
We now discuss the relationship between the agent’s the choice of garbling amount $\varepsilon$ and total welfare, $W_{\text{garb}}(p_{0},p_{1})\coloneqq\Pi_{\text{garb}}(p_{0},p_{1})+V_{\text{garb}}(p_{0},p_{1})$. For notational convenience, let $W_{\text{garb}}(\varepsilon)\coloneqq W_{\text{garb}}(p_{0}(\varepsilon),p_{1}(\varepsilon))$.
First, we show that relative to the fully concealed setting, the marginal effect of revealing any information on total welfare is initially positive relative to full concealment.
Lemma 8 (More information initially increases total welfare).
$W_{\text{garb}}^{\prime}(0)\geq 0$. The inequality is strict if $\Pi_{\text{garb}}^{\prime}(0)>0$, which is true under strict concavity of $\Pi_{0}(p),\Pi_{1}(p)$ (Assumption 1) and the MLRP (Assumption 3).
Second, we also know that the optimal amount of garbling chosen by the agent also increases total welfare over the fully concealed setting.
Lemma 9 (Optimal garbling increases welfare over concealment).
Let $\varepsilon^{*}\in\arg\max_{\varepsilon\in[0,1]}V_{\text{garb}}(\varepsilon)$. Then $W_{\text{garb}}(\varepsilon^{*})\geq W_{\text{garb}}(0)$.
Intuitively, Lemma 9 follows from the fact that the principal is never hurt by additional information. Given that the agent benefits from their optimal garbling choice, total welfare must increase.
The increase in total welfare is important from a mechanistic standpoint: consider a principal deciding whether or not to allow for agent garbling in an information transfer mechanism. There exist cases when the agent would reveal some amount of garbled information in the game, but would choose to conceal if only allowed to choose between concealment and revelation (see Figure 5 for an example). Thus, the garbling mechanism leads to higher total welfare in those situations.
While the agent’s chosen amount of garbling $\varepsilon^{*}$ improves total welfare over concealment, the question remains of where $\varepsilon^{*}$ falls relative to the optimal amount of garbling that maximizes total welfare. Lemma 10 shows that the optimal amount of garbling that maximizes total welfare must necessarily reveal at least as much information as the optimal amount of garbling chosen by the agent. Under Assumptions 1 and 3, if $\varepsilon^{*}<1$, then the $\varepsilon$ that maximizes total welfare is strictly higher than the $\varepsilon$ that maximizes agent utility.
Lemma 10 (More information increases total welfare relative to agent optimal garbling).
$W_{\text{garb}}^{\prime}(\varepsilon)\geq V_{\text{garb}}^{\prime}(\varepsilon)$ for all $\varepsilon\in[0,1]$. The inequality is strict if $\Pi_{\text{garb}}^{\prime}(0)>0$, which is true under strict concavity of $\Pi_{0}(p),\Pi_{1}(p)$ (Assumption 1) and the MLRP (Assumption 3).
4.7 Garbling vs. Restricted Price Discrimination
Our garbling model expands the agent’s action space from a binary choice between full concealment and full revelation for a given $X$ to a continuous choice of revealing a garbled version parameterized by $\varepsilon$. Thus, $\varepsilon$ interpolates between the full concealment and full revelation settings.
Garbling is not the only way to interpolate between the full concealment and full revelation settings. In the price discrimination literature, there is an established model that interpolates between full price discrimination and no price discrimination by restricting that price difference between market segments can be no greater than some parameter $r$. Wright [1993] models this restriction as arising from a “cost of transport” or arbitrage between two markets. Aguirre et al. [2010] apply this interpolation by analyzing the marginal effect of $r$ on total welfare. We refer to this interpolation using $r$ as a restricted price discrimination model. We also leveraged this technique to analyze the agent’s utility in Section 3.2.
We now discuss in detail how the garbling model that we have introduced compares with this restricted price discrimination model. Specifically, we consider how the interpolation between concealment and revelation introduced through varying $\varepsilon$ in our garbling model compares to interpolation using a constraint parameter $r$.
In fact, the trajectory of the principal and agents’ utilities as $r$ varies is different from the trajectory of the principal and agents’ utilities as $\varepsilon$ varies. Most importantly to our setting, there is a qualitative difference between the functions $V_{\text{const}}(r)$ and $V_{\text{garb}}(\varepsilon)$. Lemma 1 shows that the value of $r$ that maximizes $V_{\text{const}}(r)$ always corresponds with either full concealment or full revelation. However, under the same conditions, the value of $\varepsilon$ that maximizes $V_{\text{garb}}(\varepsilon)$ is not always at the extremes, and is often somewhere in between 00 and $1$. This is significant in our setting since the agent’s power to choose $\varepsilon$ is directly built into the game, and the existence of an optimal $\varepsilon\in(0,1)$ means that the agent benefits from the additional degree of freedom in their action space.
To visualize this difference between these interpolation methods, we can further map the combinations of principal and agent value onto the surplus triangle from Bergemann et al. [2015]. Figure 5 shows an example where there exists an intermediate value $\varepsilon$ that the agent prefers over both concealment and revelation. In summary, both this example and Lemma 1 show that while there sometimes exist intermediate values $\varepsilon$ that the agent prefers over both concealment and revelation, this is notably not true for intermediate restrictions $r$ to the amount of price discrimination.
Principal utility, $\Pi$
Agent utility, $V$
Figure 5: Trajectories of principal and agent utilities over $\varepsilon$ and $r$, mapped onto the triangle of possible combinations of principal and agent utilities from Bergemann et al. [2015]. Here, cost is distributed as a mixture of exponentials with $C|X=1\sim\text{Exp}(\frac{1}{\lambda_{1}})$, $C|X=0\sim\text{Exp}(\frac{1}{\lambda_{0}})$, with $\lambda_{0}=0.5$, $\lambda_{1}=0.01$. The point $A$ corresponds to the concealed setting ($V_{\text{con}}(p^{*}),\Pi_{\text{con}}(p^{*})$), and the point $F$ corresponds to the revealed setting ($V_{\text{rev}}(p_{0}^{*},p_{1}^{*}),\Pi_{\text{rev}}(p_{0}^{*},p_{1}^{*})$). The solid blue line shows all combinations of $\Pi_{\text{garb}}(\varepsilon),V_{\text{garb}}(\varepsilon)$ for $\varepsilon\in[0,1]$. The dashed orange line shows all combinations of $V_{\text{const}}(r),\Pi_{\text{const}}(r)$ for $r\in[0,p_{0}^{*}-p_{1}^{*}]$. First, note that the agent’s utility at $A$ is higher than at $F$, so the agent prefers concealment over revelation for this particular $X$. However, there exists a point along the $\varepsilon$ trajectory in which $V_{\text{garb}}(\varepsilon)$ achieves higher agent utility than the point $A$. However, this is not true of the $r$ trajectory. In general, Lemma 1 shows that intermediate values of $r$ will always be dominated by either the fully concealed or fully revealed settings.
5 Conclusions and Future Work
We have presented a model in which an agent decides whether or not to reveal the observability of an environmental variable to a principal. We first considered a simple action space where the agent must choose between concealing or revealing a given variable $X$, and we later relaxed this to a continuous action space where the agent can choose to reveal a garbled version of the given environmental variable $X$. In both cases, we gave sufficient conditions on the conditional cost distributions for the agent to prefer revelation over concealment, concealment over revelation, and garbling over revelation. Our agency game has a price discrimination analog, which makes it possible to leverage existing results from price discrimination to analyze total welfare; however, our model also offers a qualitatively different perspective and interpolation method via garbling.
More broadly, this work was motivated by analyzing a mechanism by which one might discover unknown unknowns. Even as data and computational methods become increasingly sophisticated and widely available, this problem of discovery of which metrics or variables to analyze continues to permeate the natural sciences, social sciences, and engineering. In this work, we’ve explored one avenue for discovery of metrics in a setting of information asymmetry where relevant environmental variables are unknown to a principal, but known to an evaluated agent. There are other possibilities for formulating the question of who holds relevant information, and when they would be willing to share it. For example, one could model incentives for third-party individuals to offer new metrics, or perhaps bi-directional information transfer where a principal and agent both hold distinct information. The key element of our work that may be worth retaining in alternative information design frameworks is the property that the variable itself may be unknown to the information receiver.
Going beyond information design, our broader motivation is to apply these insights to help design mechanisms that would benefit all players. For example, if sharing information would increase total welfare, but decrease an agent’s private utility, then perhaps a wealth transfer mechanism exists to adequately compensate the agent for revealing the information. Or, perhaps there exist auction or tournament frameworks that may further encourage information sharing. Finally, a particularly interesting avenue for future work would be to consider the possibility of cooperation and competition between multiple agents in this agency game with information transfer.
Acknowledgements
This work is supported by Google Research and the European Union (ERC-2022-SYG-OCEAN-101071601).
Appendix A Proofs from Section 3
Here we give full proofs from Section 3.
A.1 Proof of Proposition 1
To prove Proposition 1, we first reorganize the difference between the agent’s concealed and revealed utilities:
$V_{\text{con}}(p^{*})-V_{\text{rev}}(p_{1}^{*},p_{1}^{*})$ $=\theta\Delta V_{1}-(1-\theta)\Delta V_{0}$
where
$\Delta V_{0}\coloneqq V_{0}(p_{0}^{*})-V_{0}(p^{*});\quad\Delta V_{1}\coloneqq V_{1}(p^{*})-V_{1}(p_{1}^{*}).$
We begin with Lemma 11 below which upper bounds $\Delta V_{0}$.
Lemma 11.
For any concave and continuously differentiable CDF $F_{0}$, $\Delta V_{0}\leq p_{0}^{*}-\bar{p}$ for any $\bar{p}<p_{0}^{*}$.
Proof.
$\Delta V_{0}\leq p_{0}^{*}-\bar{p}$ if
$\frac{V_{0}(p_{0}^{*})-V_{0}(\bar{p})}{p_{0}^{*}-\bar{p}}\leq 1.$
We upper bound this difference by differentiating $V_{0}$:
[…]
Since $F_{0}$ is a concave and continuously differentiable CDF, by the mean value theorem,
$\frac{V_{0}(p_{0}^{*})-V_{0}(\bar{p})}{p_{0}^{*}-\bar{p}}\leq\sup_{p}V_{0}^{\prime}(p)=\sup_{p}F_{0}(p)\leq 1.$
We now leverage Lemma 11 to prove the full proposition.
Proposition 1 (Sufficient concealment condition with zero-cost type).
Suppose $F_{0}$ is a concave and continuously differentiable CDF. Suppose $C|X=1$ takes value 00 with probability $1$. Suppose the ratio $\frac{F_{0}(p)}{f_{0}(p)}$ is strictly monotone increasing for $p>0$. Then $V_{\text{con}}(p^{*})>V_{\text{rev}}(p_{0}^{*},p_{1}^{*})$ if
$\theta>(1-\theta)\frac{1}{\eta((1-\theta)p_{0}^{*})}-\frac{1}{\eta_{0}\left(p_{0}^{*}\right)}$
where $\eta(p)=\frac{p(1-\theta)f_{0}(p)}{(1-\theta)F_{0}(p)+\theta}$ and $\eta_{0}(p)=\frac{pf_{0}(p)}{F_{0}(p)}$ are the respective price elasticities for task completion quantity for the mixture distribution $C$ and the conditional distribution $C|X=0$.
Proof.
We consider the extreme case where $C|X=1$ has value 00 with probability $1$. For this distribution of $C|X=1$, we show that Equation (3) implies that
$\theta\Delta V_{1}>(1-\theta)\Delta V_{0}.$ (13)
First, for any nonzero $C|X=0$, we have that $p^{*}<p_{0}^{*}$. Thus, Lemma 11 gives that
$(1-\theta)\Delta V_{0}\leq(1-\theta)(p_{0}^{*}-p^{*}).$
Next, we further upper bound this by showing that for any $F_{0}$ that satisfies Equation (3),
$(1-\theta)(p_{0}^{*}-p^{*})<\theta p^{*}.$ (14)
Equation (13) then follows from the fact that $\Delta V_{1}=p^{*}$ when $C|X=1$ is always zero.
We now prove that equation (14) holds under equation (3). First, note that
$(1-\theta)(p_{0}^{*}-p^{*})<\theta p^{*}\iff(1-\theta)p_{0}^{*}<p^{*}.$
Since $F_{0}$ is concave and continuously differentiable, $p^{*}$ satisfies the following first-order condition:
$p^{*}+\frac{(1-\theta)F_{0}(p^{*})+\theta}{(1-\theta)f_{0}(p^{*})}=b.$ (15)
Since $\frac{F_{0}(p)}{f_{0}(p)}$ is strictly monotone increasing for $p>0$ and $F_{0}(p)$ is concave, $p+\frac{(1-\theta)F_{0}(p)+\theta}{(1-\theta)f_{0}(p)}$ is also strictly monotone increasing for $p>0$. Therefore, $(1-\theta)p_{0}^{*}<p^{*}$ if and only if
$(1-\theta)p_{0}^{*}+\frac{(1-\theta)F_{0}((1-\theta)p_{0}^{*})+\theta}{(1-\theta)f_{0}((1-\theta)p_{0}^{*})}<b.$
We also have that $p_{0}^{*}$ satisfies the first-order condition
$p_{0}^{*}+\frac{F_{0}(p_{0}^{*})}{f_{0}(p_{0}^{*})}=b.$
Therefore, $(1-\theta)p_{0}^{*}<p^{*}$ if and only if
$(1-\theta)p_{0}^{*}+\frac{(1-\theta)F_{0}((1-\theta)p_{0}^{*})+\theta}{(1-\theta)f_{0}((1-\theta)p_{0}^{*})}<p_{0}^{*}+\frac{F_{0}(p_{0}^{*})}{f_{0}(p_{0}^{*})},$
which is equivalent to the condition in equation (3).
A.2 Proofs for Propositions 2 and 3
Here we give proofs for Propositions 2 and 3. These analyses parallel those of Aguirre et al. [2010] for the effects of price discrimination on total welfare.
We first prove Lemma 1, which follows from Assumption 1 and 2.
Lemma 1.
Under Assumptions 1 and 2, $V_{\text{const}}(r)$ is strictly quasi-convex for $r\in[0,p_{0}^{*}-p_{1}^{*}]$. That is, if there exists $\hat{r}\in[0,p_{0}^{*}-p_{1}^{*}]$ such that $V_{\text{const}}(\hat{r})=0$, then $V_{\text{const}}^{\prime\prime}(\hat{r})>0$.
Proof.
The constraint in equation (4) is binding when $r\in[0,p_{0}^{*}-p_{1}^{*}]$. Therefore, the optimization problem in equation (4) can be rewritten as
$\max_{p_{1}}\;\;\Pi_{1}(p_{1})+\Pi_{0}(p_{1}+r),$
yielding a first-order condition that $\Pi_{1}^{\prime}(p_{1})+\Pi_{0}^{\prime}(p_{1}+r)=0$. Further differentiating this first-order condition, as done by Aguirre et al. [2010], yields that
$p_{1}^{\prime}(r)=\frac{-\Pi_{0}^{\prime\prime}(p_{0}(r))}{\Pi_{0}^{\prime\prime}(p_{0}(r))+\Pi_{1}^{\prime\prime}(p_{1}(r))}.$
A similar method shows that
$p_{0}^{\prime}(r)=\frac{\Pi_{1}^{\prime\prime}(p_{1}(r))}{\Pi_{0}^{\prime\prime}(p_{0}(r))+\Pi_{1}^{\prime\prime}(p_{1}(r))}.$
Thus, we have that
$\begin{split}V_{\text{const}}^{\prime}(r)&=(1-\theta)V_{0}(p_{0}(r))p_{0}^{\prime}(r)+\theta V_{1}(p_{1}(r))p_{1}^{\prime}(r)\\ &=\left(\frac{-\Pi_{1}^{\prime\prime}(p_{1}(r)\Pi_{0}^{\prime\prime}(p_{0}(r)))}{\Pi_{0}^{\prime\prime}(p_{0}(r))+\Pi_{1}^{\prime\prime}(p_{1}(r))}\right)\left(\frac{\theta V_{1}^{\prime}(p_{1}(r))}{\Pi_{1}^{\prime\prime}(p_{1}(r))}-\frac{(1-\theta)V_{0}^{\prime}(p_{0}(r))}{\Pi_{0}^{\prime\prime}(p_{0}(r))}\right)\\ &=\left(\frac{-\Pi_{1}^{\prime\prime}(p_{1}(r)\Pi_{0}^{\prime\prime}(p_{0}(r)))}{\Pi_{0}^{\prime\prime}(p_{0}(r))+\Pi_{1}^{\prime\prime}(p_{1}(r))}\right)\left(\theta w_{1}(p_{1}(r))-(1-\theta)w_{0}(p_{0}(r))\right)\\ \end{split}$ (16)
where $w_{x}(p)\coloneqq\frac{V_{x}^{\prime}(p)}{\Pi_{x}^{\prime\prime}(p)}$.
Taking the second derivative, we have that
$V_{\text{const}}^{\prime\prime}(r)$ $=\left(\frac{-\Pi_{1}^{\prime\prime}(p_{1}(r)\Pi_{0}^{\prime\prime}(p_{0}(r)))}{\Pi_{0}^{\prime\prime}(p_{0}(r))+\Pi_{1}^{\prime\prime}(p_{1}(r))}\right)\left(\theta w_{1}^{\prime}(p_{1}(r))p_{1}^{\prime}(r)-(1-\theta)w_{0}^{\prime}(p_{0}(r))p_{0}^{\prime}(r)\right)$
$\quad+(\theta w_{1}(p_{1}(r))-(1-\theta)w_{0}(p_{0}(r)))\frac{\partial}{\partial r}\left(\frac{-\Pi_{1}^{\prime\prime}(p_{1}(r)\Pi_{0}^{\prime\prime}(p_{0}(r)))}{\Pi_{0}^{\prime\prime}(p_{0}(r))+\Pi_{1}^{\prime\prime}(p_{1}(r))}\right).$
The first term $\left(\frac{-\Pi_{1}^{\prime\prime}(p_{1}(r)\Pi_{0}^{\prime\prime}(p_{0}(r)))}{\Pi_{0}^{\prime\prime}(p_{0}(r))+\Pi_{1}^{\prime\prime}(p_{1}(r))}\right)$ is positive by strict concavity given by Assumption 1.
If $V_{\text{const}}^{\prime}(\hat{r})=0$, then $\theta w_{1}(p_{1}(\hat{r}))-(1-\theta)w_{0}(p_{0}(\hat{r}))=0$. By the DRC, $w_{1}^{\prime}(p_{1}(\hat{r}))p_{1}^{\prime}(\hat{r})>0$ since $w_{1}^{\prime}(p_{1}(\hat{r}))<0$ and $p_{1}^{\prime}(\hat{r})<0$. Similarly, $w_{0}^{\prime}(p_{0}(\hat{r}))p_{0}^{\prime}(\hat{r})<0$. Therefore, $V_{\text{const}}^{\prime\prime}(\hat{r})>0$.
Given Lemma 1, we now prove Propositions 2 and 3 by signing the derivative $V_{\text{const}}^{\prime}(r)$ for extreme values of $r$.
Proposition 2 (Sufficient concealment condition under DRC).
Under Assumptions 1 and 2, $V_{\text{con}}(p^{*})>V_{\text{rev}}(p_{0}^{*},p_{1}^{*})$ if
$\frac{(1-\theta)(b-p_{0}^{*})}{2-\sigma_{0}(p_{0}^{*})}<\frac{\theta(b-p_{1}^{*})}{2-\sigma_{1}(p_{1}^{*})},$
where $\sigma_{x}(p)=\frac{F_{x}(p)f^{\prime}_{x}(p)}{f_{x}^{2}(p)}$ is the curvature of the inverse of the task completion quantity function $F_{x}(p)$.
Proof.
If $V_{\text{const}}(r)$ is strictly monotone decreasing in $r$, then $V_{\text{con}}(p^{*})>V_{\text{rev}}(p_{0}^{*},p_{1}^{*})$. Since $V_{\text{const}}(r)$ is strictly quasi-convex, a sufficient condition for $V_{\text{const}}(r)$ to be strictly monotone decreasing is $V_{\text{const}}^{\prime}(p_{0}^{*}-p_{1}^{*})<0$.
By equation (16), we have that $V_{\text{const}}^{\prime}(p_{0}^{*}-p_{1}^{*})<0$ if $\theta w_{1}(p_{1}^{*})-(1-\theta)w_{0}(p_{0}^{*})<0$. By the first-order condition that
$b-p_{x}^{*}=\frac{F_{x}(p_{x}^{*})}{f_{x}(p_{x}^{*})},$ (17)
we have that
$w_{x}(p_{x}^{*})=\frac{F_{x}(p_{x}^{*})}{\Pi^{\prime\prime}_{x}(p_{x}^{*})}=\frac{b-p_{x}^{*}}{\Pi^{\prime\prime}_{x}(p_{x}^{*})/f_{x}(p_{x}^{*})}$
Note that
$\Pi^{\prime\prime}_{x}(p)=-2f_{x}(p)+f_{x}^{\prime}(p)(b-p)$
Therefore, also applying the first-order condition from equation (17), we have
$\Pi^{\prime\prime}_{x}(p_{x}^{*})/f_{x}(p_{x}^{*})=-2+\frac{F_{x}(p)f^{\prime}_{x}(p)}{f_{x}^{2}(p)}=-2+\sigma_{x}(p).$
A similar argument yields Proposition 3.
Proposition 3 (Sufficient revelation condition under DRC).
Under Assumptions 1 and 2, $V_{\text{con}}(p^{*})<V_{\text{rev}}(p_{0}^{*},p_{1}^{*})$ if
$\frac{2+L(p^{*})\alpha_{1}(p^{*})}{2+L(p^{*})\alpha_{0}(p^{*})}>\frac{\theta F_{1}(p^{*})/f_{1}(p^{*})}{(1-\theta)F_{0}(p^{*})/f_{0}(p^{*})},$
where $L(p)=\frac{b-p}{p}$ is the Lerner index, and $\alpha_{x}(p)=\frac{-pf_{x}^{\prime}(p)}{f_{x}(p)}$ is the curvature of the task completion quantity function $F_{x}(p)$.
Proof.
If $V_{\text{const}}(r)$ is strictly monotone increasing in $r$, then $V_{\text{con}}(p^{*})<V_{\text{rev}}(p_{0}^{*},p_{1}^{*})$. Since $V_{\text{const}}(r)$ is strictly quasi-convex, a sufficient condition for $V_{\text{const}}(r)$ to be strictly monotone increasing is $V_{\text{const}}^{\prime}(0)>0$.
By equation (16), we have that $V_{\text{const}}^{\prime}(0)>0$ if $\theta w_{1}(p^{*})-(1-\theta)w_{0}(p^{*})>0$.
$w_{x}(p^{*})=\frac{F_{x}(p^{*})/f_{x}(p^{*})}{-2+(b-p^{*})(f_{x}^{\prime}(p^{*})/f_{x}(p^{*}))}=\frac{F_{x}(p^{*})/f_{x}(p^{*})}{-2-L(p^{*})\alpha_{x}(p^{*})}$
Therefore, $\theta w_{1}(p^{*})-(1-\theta)w_{0}(p^{*})>0$ if
$\frac{\theta F_{1}(p^{*})/f_{1}(p^{*})}{2+L(p^{*})\alpha_{1}(p^{*})}<\frac{(1-\theta)F_{0}(p^{*})/f_{0}(p^{*})}{2+L(p^{*})\alpha_{0}(p^{*})}$
A.3 Comparison of decreasing ratio condition to increasing ratio condition
The results in Section 3.2.2 depend on the decreasing ratio condition (DRC) given in Assumption 2. This is analogous to the “increasing ratio condition (IRC)” from Aguirre et al. [2010], which says that the ratio $\frac{W_{x}^{\prime}(p)}{\Pi_{x}^{\prime\prime}(p)}$ is increasing in $p$. We now discuss in more detail the relationship between the DRC and the IRC, including sufficient conditions under which the DRC holds. In introducing the IRC, Aguirre et al. [2010] describe a “very large set of demand functions” for which the IRC holds. Aguirre et al. [2010] give sufficient conditions for the IRC to hold in Appendix B from their paper, which includes linear functions and exponential and constant elasticity functions.
For all of the sufficient conditions that Aguirre et al. [2010] proposes for the IRC, these are also sufficient conditions for the DRC if paired with the additional condition that $\frac{F_{x}(p)}{f_{x}(p)}$ is increasing in $p$ for $x\in\{0,1\}$. For example, a specific sufficient condition for the DRC, which also implies the IRC, is the following: Let $\sigma(p)=\frac{F(p)f^{\prime}(p)}{f(p)^{2}}$. If $\sigma(p)\leq 1$, and $\alpha(p)=-\frac{pf^{\prime}(p)}{f(p)}$ is non-decreasing and positive in $p$, then the DRC holds. The IRC would also hold. A similar analogy can be made for all other conditions given in Appendix B of Aguirre et al. [2010].
A.4 Proofs from Section 3.4
Lemmas 2 and 3 show that the principal always benefits from more information being revealed. Assumption 3 further implies that the principal strictly benefits from revelation.
Lemma 2 (Principal prefers revelation).
Revealing $X$ never decreases the value of the principal: $\Pi_{\text{rev}}(\rho^{*})\geq\Pi_{\text{con}}(p^{*})$, where $\rho^{*}\in\arg\max_{\rho}\Pi_{\text{rev}}(\rho)$. Revealing $X$ strictly increases the value of the principal only if $X$ and $C$ are not independent.
Proof.
Assuming that the principal’s feasible set of payments does not change between markets, the solution $\hat{\rho}(x)=p^{*}$ is in the feasible set of the principal’s optimization problem with information revealed. Therefore,
$\max_{\rho}\Pi_{\text{rev}}(\rho)\geq\Pi_{\text{rev}}(\hat{\rho})=\Pi_{\text{con}}(p^{*}).$
If $X$ and $C$ are independent, we have $F_{x}=F$ for all $x\in\mathcal{X}$, so
$\max_{\rho}\Pi_{\text{rev}}(\rho)=\max_{\rho}\mathbb{E}[F(\rho(X))(b-\rho(X))]$
By Jensen’s inequality, we have
$\max_{\rho}\mathbb{E}[F(\rho(X))(b-\rho(X))]\leq\mathbb{E}[\max_{\rho}F(\rho(X))(b-\rho(X))]=\mathbb{E}[F(p^{*})(b-p^{*})]=\Pi_{\text{con}}(p^{*}).$
Therefore, if $X$ and $C$ are independent, then $\Pi_{\text{rev}}(\rho^{*})\leq\Pi_{\text{con}}(p^{*})$. ∎
Lemma 3 (Principal strictly benefits from revelation).
Let $F_{0}$ and $F_{1}$ be continuously differentiable CDFs. If the MLRP holds (Assumption 3), then the principal strictly benefits when $X$ is revealed: $\Pi_{\text{rev}}(p_{0}^{*},p_{1}^{*})>\Pi_{\text{con}}(p^{*})$.
Proof.
Since $F_{0}$, $F_{1}$ are continuously differentiable, the first-order necessary conditions hold for the optimal payments $p_{0}^{*}$, $p_{1}^{*}$ in equation (17). By these conditions, $p_{0}^{*}=p_{1}^{*}$ only if there exists a value $p$ such that
$p+\frac{F_{0}(p)}{f_{0}(p)}=p+\frac{F_{1}(p)}{f_{1}(p)}=b.$
Such a value $p$ cannot exist if $\frac{F_{0}(p)}{f_{0}(p)}\neq\frac{F_{1}(p)}{f_{1}(p)}$ for all $p$. The MLRP implies that $\frac{F_{0}(p)}{f_{0}(p)}<\frac{F_{1}(p)}{f_{1}(p)}$ for all $p$; therefore, $p_{0}^{*}\neq p_{1}^{*}$, and maximum value for $\Pi_{\text{rev}}$ is strictly greater than the maximum value for $\Pi_{\text{con}}$. ∎
A.5 Proofs from Section 3.5
Lemma 4.
Total welfare increases under revelation ($W_{\text{rev}}(p_{0}^{*},p_{1}^{*})-W_{\text{con}}(p^{*})$) only if task completion quantity increases under revelation,
$F(p^{*})<(1-\theta)F_{0}(p_{0}^{*})+\theta F_{1}(p_{1}^{*}).$
Proof.
[…]
Upper bounding this:
[…]
The right hand term is exactly the difference in output:
[…]
$=(1-\theta)F_{0}(p_{0}^{*})+\theta F_{1}(p_{1}^{*})-((1-\theta)F_{0}(p^{*})+\theta F_{1}(p^{*}))$
$=(1-\theta)F_{0}(p_{0}^{*})+\theta F_{1}(p_{1}^{*})-F(p^{*})$
$\implies W_{\text{rev}}(p_{0}^{*},p_{1}^{*})-W_{\text{con}}(p^{*})\leq(b-p^{*})\left((1-\theta)F_{0}(p_{0}^{*})+\theta F_{1}(p_{1}^{*})-F(p^{*})\right)$
Since $(b-p^{*})>0$, if $(1-\theta)F_{0}(p_{0}^{*})+\theta F_{1}(p_{1}^{*})-F(p^{*})\leq 0$, then $W_{\text{rev}}(p_{0}^{*},p_{1}^{*})-W_{\text{con}}(p^{*})\leq 0$. ∎
Appendix B Proofs from Section 4
Here we give proofs for results for the garbling model presented in Section 4.
B.1 Proofs from Section 4.2
Lemma 5 (Monotonic price changes).
Suppose $\gamma=\theta=\frac{1}{2}$. Suppose $F_{0},F_{1}$ are continuously differentiable CDFs, the principal’s utility is strictly concave (Assumption 1), and the MLRP holds (Assumption 3). Then $p_{0}^{\prime}(\varepsilon)>0$ and $p_{1}^{\prime}(\varepsilon)<0$ for all $\varepsilon\in[0,1]$.
Proof.
$p_{0}(\varepsilon),p_{1}(\varepsilon)$ must satisfy first-order necessary conditions for optimality:
$p_{0}(\varepsilon)+\frac{\frac{1+\varepsilon}{2}F_{0}(p_{0}(\varepsilon))+\frac{1-\varepsilon}{2}F_{1}(p_{0}(\varepsilon))}{\frac{1+\varepsilon}{2}f_{0}(p_{0}(\varepsilon))+\frac{1-\varepsilon}{2}f_{1}(p_{0}(\varepsilon))}=b;\quad p_{1}(\varepsilon)+\frac{\frac{1+\varepsilon}{2}F_{1}(p_{1}(\varepsilon))+\frac{1-\varepsilon}{2}F_{0}(p_{1}(\varepsilon))}{\frac{1+\varepsilon}{2}f_{1}(p_{1}(\varepsilon))+\frac{1-\varepsilon}{2}f_{0}(p_{1}(\varepsilon))}=b$
Differentiating these first-order conditions, we have:
$p_{0}^{\prime}(\varepsilon)=\frac{\Pi_{0}^{\prime}(p_{0}(\varepsilon))-\Pi_{1}^{\prime}(p_{0}(\varepsilon))}{-2\Pi_{Y=0}^{{}^{\prime\prime}}(p_{0}(\varepsilon))};\quad p_{1}^{\prime}(\varepsilon)=\frac{\Pi_{1}^{\prime}(p_{1}(\varepsilon))-\Pi_{0}^{\prime}(p_{1}(\varepsilon))}{-2\Pi_{Y=1}^{{}^{\prime\prime}}(p_{1}(\varepsilon))},$ (18)
where
$\Pi_{Y=0}(p)=\frac{1+\varepsilon}{2}\Pi_{0}(p)+\frac{1-\varepsilon}{2}\Pi_{1}(p);\quad\Pi_{Y=1}(p)=\frac{1+\varepsilon}{2}\Pi_{1}(p)+\frac{1-\varepsilon}{2}\Pi_{0}(p).$
Strict concavity from Assumption 1 makes both denominators of $p_{x}^{\prime}(\varepsilon)$ positive.
The MLRP also implies that $p_{0}(\varepsilon)<p_{0}^{*}$ and $p_{1}(\varepsilon)>p_{1}^{*}$ for any $\varepsilon$. Therefore, by strict concavity of $\Pi_{x}(p)$, we have $\Pi_{0}^{\prime}(p_{0}(\varepsilon))>0$, and $\Pi_{1}^{\prime}(p_{0}(\varepsilon))<0$, implying that $p_{0}^{\prime}(\varepsilon)>0$. Similarly, $\Pi_{1}^{\prime}(p_{1}(\varepsilon))<0$, and $\Pi_{0}^{\prime}(p_{1}(\varepsilon))>0$, implying that $p_{1}^{\prime}(\varepsilon)<0$. ∎
B.2 Proofs from Section 4.3
We first expand $V_{\text{garb}}(\varepsilon)$ for $\gamma=\theta=\frac{1}{2}$.
$V_{\text{garb}}(\varepsilon)=\frac{1}{2}\left(\frac{1+\varepsilon}{2}V_{0}(p_{0}(\varepsilon))+\frac{1-\varepsilon}{2}V_{1}(p_{0}(\varepsilon))+\frac{1+\varepsilon}{2}V_{1}(p_{1}(\varepsilon))+\frac{1-\varepsilon}{2}V_{0}(p_{1}(\varepsilon))\right).$ (19)
To prove Proposition 4, we first give Lemma 12 to handle the anchored zero-cost agent.
Lemma 12.
Suppose $C|X=1$ takes value 00 with probability $1$. Suppose $f_{0}$ is bounded: $f_{0}(p)<B$ for all $p$ in the support. Then for any fixed value $b>0$, there exists $\delta>0$ such that for any $\varepsilon>\delta$, $p_{1}(\varepsilon)=0$.
Proof.
Define $h(p,\varepsilon)=p+\frac{1}{f_{0}(p)}\frac{1+\varepsilon}{1-\varepsilon}+\frac{F_{0}(p)}{f_{0}(p)}$. By choosing $\delta$ close to 1, we can make the term $\frac{1+\delta}{1-\delta}$ arbitrarily large, and consequently $\frac{1}{f_{0}(p_{1})}\frac{1+\delta}{1-\delta}$ arbitrarily large, since $f_{0}(p)>0$. Then for any $b$, we choose $\delta$ close enough to 1 such that $\frac{1}{B}\frac{1+\delta}{1-\delta}>b$. ∎
Proposition 4 (Sufficient garbling condition with zero-cost type).
Suppose $\gamma=\theta=\frac{1}{2}$. Suppose $C|X=1$ takes value 0 with probability $1$. Suppose $F_{0}$ is continuously differentiable and $f_{0}(c)$ is bounded. $V_{\text{garb}}(\varepsilon)$ is maximized at $\varepsilon^{*}<1$ if
$\frac{v-p_{0}^{*}}{2-\sigma_{0}(p_{0}^{*})}<g_{0}(p_{0}^{*}),$
[…]
Proof.
We show that the condition in equation (11) implies that $V_{\text{garb}}^{\prime}(1)<0$.
For $C|X=1$ taking value 00 with probability $1$, we have $V_{1}(p)=p$. Substituting this into equation (19),
$V_{\text{garb}}(\varepsilon)=\frac{1}{2}\left(\frac{1+\varepsilon}{2}V_{0}(p_{0}(\varepsilon))+\frac{1-\varepsilon}{2}p_{0}(\varepsilon)+\frac{1+\varepsilon}{2}p_{1}(\varepsilon)+\frac{1-\varepsilon}{2}V_{0}(p_{1}(\varepsilon))\right).$
Differentiating this, we have
$2V_{\text{garb}}^{\prime}(\varepsilon)=$ $p_{0}^{\prime}(\varepsilon)\left(\frac{1+\varepsilon}{2}F_{0}(p_{0}(\varepsilon))+\frac{1-\varepsilon}{2}\right)+p_{1}^{\prime}(\varepsilon)\left(\frac{1+\varepsilon}{2}+\frac{1-\varepsilon}{2}F_{0}(p_{1}(\varepsilon))\right)$
$+\frac{1}{2}\left((V_{0}(p_{0}(\varepsilon))-V_{0}(p_{1}(\varepsilon)))-(p_{0}(\varepsilon)-p_{1}(\varepsilon))\right)$
Evaluating this derivative at $\varepsilon=1$, Lemma 12 implies that $p_{1}(\varepsilon)=0$ and $p_{1}^{\prime}(1)=0$.
$2V_{\text{garb}}^{\prime}(1)=p_{0}^{\prime}(1)F_{0}(p_{0}^{*})+\frac{1}{2}(V_{0}(p_{0}^{*})-p_{0}^{*}).$
$V_{0}(p_{0}^{*})-p_{0}^{*}=E[(p_{0}^{*}-C)\operatorname{\mathbbm{1}}(C<p_{0}^{*})|X=0]-p_{0}^{*}=-\frac{1}{2}g_{0}(p_{0}^{*}).$
$\implies 4V_{\text{garb}}^{\prime}(1)$ $=-g_{0}(p_{0}^{*})+2F_{0}(p_{0}^{*})p_{0}^{\prime}(1)$
Simplifying $2F_{0}(p_{0}^{*})p_{0}^{\prime}(1)$:
$2F_{0}(p_{0}^{*})p_{0}^{\prime}(1)$ $=\frac{F_{0}(p_{0}^{*})f_{0}(p_{0}^{*})}{2f_{0}(p_{0}^{*})^{2}-F_{0}(p_{0}^{*})f_{0}^{\prime}(p_{0}^{*})}$
$=\frac{\frac{F_{0}(p_{0}^{*})}{f_{0}(p_{0}^{*})}}{2-\frac{F_{0}(p_{0}^{*})f_{0}^{\prime}(p_{0}^{*})}{f_{0}(p_{0}^{*})^{2}}}$
$=\frac{v-p_{0}^{*}}{2-\sigma_{0}(p_{0}^{*})}$
Therefore,
$V_{\text{garb}}^{\prime}(1)<0\iff-g_{0}(p_{0}^{*})+\frac{v-p_{0}^{*}}{2-\sigma_{0}(p_{0}^{*})}<0.$
Proposition 5 (Sufficient garbling condition).
Suppose $\gamma=\theta=\frac{1}{2}$. Suppose $F_{0},F_{1}$ are continuously differentiable. $V_{\text{garb}}(\varepsilon)$ is maximized at $\varepsilon^{*}<1$ if
$-\Pi_{1}^{\prime}(p_{0}^{*})\left(\frac{b-p_{0}^{*}}{2-\sigma_{0}(p_{0}^{*})}\right)-\Pi_{0}^{\prime}(p_{1}^{*})\left(\frac{b-p_{1}^{*}}{2-\sigma_{1}(p_{1}^{*})}\right)<\Delta(p_{0}^{*},p_{1}^{*}).$
Proof.
Differentiating with respect to $\varepsilon$, we have
$2V_{\text{garb}}^{\prime}(\varepsilon)=$ $p_{0}^{\prime}(\varepsilon)\left(\frac{1+\varepsilon}{2}F_{0}(p_{0}(\varepsilon))+\frac{1-\varepsilon}{2}F_{1}(p_{0}(\varepsilon))\right)+p_{1}^{\prime}(\varepsilon)\left(\frac{1+\varepsilon}{2}F_{1}(p_{1}(\varepsilon))+\frac{1-\varepsilon}{2}F_{0}(p_{1}(\varepsilon))\right)$
$+\frac{1}{2}\left((V_{0}(p_{0}(\varepsilon))-V_{0}(p_{1}(\varepsilon)))-(V_{1}(p_{0}(\varepsilon))-V_{1}(p_{1}(\varepsilon)))\right)$
Substituting in the price derivatives from equation (18) and the agent utility dominance identity from Definition 1,
$2V_{\text{garb}}^{\prime}(\varepsilon)=$ $\left(\frac{\Pi_{0}^{\prime}(p_{0}(\varepsilon))-\Pi_{1}^{\prime}(p_{0}(\varepsilon))}{-2\Pi_{Y=0}^{{}^{\prime\prime}}(p_{0}(\varepsilon))}\right)\left(\frac{1+\varepsilon}{2}F_{0}(p_{0}(\varepsilon))+\frac{1-\varepsilon}{2}F_{1}(p_{0}(\varepsilon))\right)$
$+\left(\frac{\Pi_{1}^{\prime}(p_{1}(\varepsilon))-\Pi_{0}^{\prime}(p_{1}(\varepsilon))}{-2\Pi_{Y=1}^{{}^{\prime\prime}}(p_{1}(\varepsilon))}\right)\left(\frac{1+\varepsilon}{2}F_{1}(p_{1}(\varepsilon))+\frac{1-\varepsilon}{2}F_{0}(p_{1}(\varepsilon))\right)$
$+\frac{1}{2}\left(-\Delta(p_{0}(\varepsilon),p_{1}(\varepsilon))\right)$
Let
$z_{0}^{\varepsilon}(p)=\frac{V^{\prime}_{Y=0}(p)}{\Pi_{Y=0}^{{}^{\prime\prime}}(p)}=\frac{\frac{1+\varepsilon}{2}F_{0}(p)+\frac{1-\varepsilon}{2}F_{1}(p)}{\Pi_{Y=0}^{{}^{\prime\prime}}(p)},$
$z_{1}^{\varepsilon}(p)=\frac{V^{\prime}_{Y=1}(p)}{\Pi_{Y=1}^{{}^{\prime\prime}}(p)}=\frac{\frac{1+\varepsilon}{2}F_{1}(p)+\frac{1-\varepsilon}{2}F_{0}(p)}{\Pi_{Y=1}^{{}^{\prime\prime}}(p)}.$
Then
$4V_{\text{garb}}^{\prime}(\varepsilon)=$ $\left(\Pi_{0}^{\prime}(p_{0}(\varepsilon))-\Pi_{1}^{\prime}(p_{0}(\varepsilon))\right)(-z_{0}^{\varepsilon}(p_{0}(\varepsilon)))+\left(\Pi_{1}^{\prime}(p_{1}(\varepsilon))-\Pi_{0}^{\prime}(p_{1}(\varepsilon))\right)(-z_{1}^{\varepsilon}(p_{1}(\varepsilon)))$
$-\Delta(p_{0}(\varepsilon),p_{1}(\varepsilon))$
Evaluating this at $\varepsilon=1$, we have
$4V_{\text{garb}}^{\prime}(1)=$ $\left(\Pi_{0}^{\prime}(p_{0}^{*})-\Pi_{1}^{\prime}(p_{0}^{*})\right)(-z_{0}(p_{0}^{*}))+\left(\Pi_{1}^{\prime}(p_{1}^{*})-\Pi_{0}^{\prime}(p_{1}^{*})\right)(-z_{1}(p_{1}^{*}))$
$-\Delta(p_{0}^{*},p_{1}^{*})$
where
$z_{x}(p)=\frac{V^{\prime}_{x}(p)}{\Pi_{x}^{{}^{\prime\prime}}(p)}$
Note that $V^{\prime}_{x}(p_{x}^{*})=W^{\prime}_{x}(p_{x}^{*})$, so at $p^{*}$, this is identical to the $z$ function from Aguirre et al. [2010]. By first-order optimality conditions,
$-z_{x}(p_{x}^{*})=\frac{b-p_{x}^{*}}{2-\sigma_{x}(p_{x}^{*})}$
Therefore, $V_{\text{garb}}^{\prime}(1)<0$ if
$\left(-\Pi_{1}^{\prime}(p_{0}^{*})\right)\left(\frac{b-p_{0}^{*}}{2-\sigma_{0}(p_{0}^{*})}\right)+\left(-\Pi_{0}^{\prime}(p_{1}^{*})\right)\left(\frac{b-p_{1}^{*}}{2-\sigma_{1}(p_{1}^{*})}\right)<\Delta(p_{0}^{*},p_{1}^{*}).$
B.3 Proofs from Section 4.5
Lemma 6.
$\Pi_{\text{con}}(p^{*})\leq\Pi_{\text{garb}}(p_{0}(\varepsilon),p_{1}(\varepsilon))\leq\Pi_{\text{rev}}(p_{0}^{*},p_{1}^{*})$ for all $\varepsilon$.
Proof.
For $\varepsilon\in\{0,1\}$, the inequalities clearly hold. Fix $\varepsilon\in(0,1)$. For the lower bound,
$\Pi_{\text{garb}}(p_{0}(\varepsilon),p_{1}(\varepsilon))\geq\Pi_{\text{garb}}(p^{*},p^{*})=\Pi_{\text{con}}(p^{*}).$
For the upper bound, since the noise $\xi$ is independent of $X$ and $C$,
$\Pi_{\text{garb}}(p_{0},p_{1})=\phi(\theta,\gamma)\Pi_{\text{rev}}(p_{0},p_{1})+\psi(\theta,\gamma)\Pi_{\text{rev}}(p_{1},p_{0})\leq\Pi_{\text{rev}}(p_{0}^{*},p_{1}^{*})$
where $\phi,\psi$ are some positive functions of $\theta,\gamma$. ∎
Lemma 7 (Strict principal improvement).
Suppose $\gamma=\theta=\frac{1}{2}$. Suppose the principal’s utility is strictly concave (Assumption 1). If the MLRP holds (Assumption 3), then $\Pi_{\text{garb}}^{\prime}(\varepsilon)>0$ for all $\varepsilon\in[0,1]$.
Proof.
$\Pi_{\text{garb}}(\varepsilon)$ $=\frac{1}{2}(b-p_{0}(\varepsilon))\left(\frac{1+\varepsilon}{2}F_{0}(p_{0}(\varepsilon))+\frac{1-\varepsilon}{2}F_{1}(p_{0}(\varepsilon))\right)$
$+\frac{1}{2}(b-p_{1}(\varepsilon))\left(\frac{1+\varepsilon}{2}F_{1}(p_{1}(\varepsilon))+\frac{1-\varepsilon}{2}F_{0}(p_{1}(\varepsilon))\right)$
Differentiating with respect to $\varepsilon$:
$4\Pi_{\text{garb}}^{\prime}(\varepsilon)=$ $F_{0}(p_{0}(\varepsilon))(b-p_{0}(\varepsilon))-F_{0}(p_{1}(\varepsilon))(b-p_{1}(\varepsilon))$
$+F_{1}(p_{1}(\varepsilon))(b-p_{1}(\varepsilon))-F_{1}(p_{0}(\varepsilon))(b-p_{0}(\varepsilon))$
By the MLRP and strict concavity of $\Pi_{0}(p),\Pi_{1}(p)$, we have that $p_{1}(\varepsilon)<p_{0}(\varepsilon)<p_{0}^{*}$.
Strict concavity of $\Pi_{0}(p)$ and the optimality of $p_{0}^{*}$ for $\Pi_{1}$ then implies that
$F_{0}(p_{1}(\varepsilon))(v-p_{1}(\varepsilon))<F_{0}(p_{0}(\varepsilon))(v-p_{0}(\varepsilon)).$
Similarly, by the MLRP and strict concavity of $\Pi_{0}(p),\Pi_{1}(p)$, we have that $p_{1}^{*}<p_{1}(\varepsilon)<p_{0}(\varepsilon)$. Strict concavity of $\Pi_{1}(p)$ and the optimality of $p_{1}^{*}$ for $\Pi_{1}$ implies that
$F_{1}(p_{0}(\varepsilon))(v-p_{0}(\varepsilon))<F_{1}(p_{1}(\varepsilon))(v-p_{1}(\varepsilon)).$
B.4 Proofs from Section 4.6
Lemma 8 (Reducing garbling initially increases total welfare).
Relative to full concealment with $\varepsilon=0$, revealing $Y$ with some garbled noise initially does not decrease total welfare: $W_{\text{garb}}^{\prime}(0)\geq 0$. The inequality is strict if $\Pi_{\text{garb}}^{\prime}(0)>0$, which is true under strict concavity of $\Pi_{0}(p),\Pi_{1}(p)$ (Assumption 1) and the MLRP (Assumption 3).
Proof.
$W_{\text{garb}}^{\prime}(\varepsilon)=V_{\text{garb}}^{\prime}(\varepsilon)+\Pi_{\text{garb}}^{\prime}(\varepsilon)$
$V_{\text{garb}}^{\prime}(0)=0$, so $W_{\text{garb}}^{\prime}(0)\geq 0$. The strict inequality comes from applying Lemma 7 that $\Pi_{\text{garb}}^{\prime}(0)>0$. ∎
Lemma 9 (Optimal garbling increases welfare over concealment).
Let $\varepsilon^{*}\in\arg\max_{\varepsilon\in[0,1]}V_{\text{garb}}(\varepsilon)$. Then $W_{\text{garb}}(\varepsilon^{*})\geq W_{\text{garb}}(0)$.
Proof.
$W_{\text{garb}}(\varepsilon^{*})=V_{\text{garb}}(\varepsilon^{*})+\Pi_{\text{garb}}(\varepsilon^{*})$
Lemma 6 implies $\Pi_{\text{garb}}(\varepsilon^{*})\geq\Pi_{\text{garb}}(0)$. By optimality of $\varepsilon^{*}$, $V_{\text{garb}}(\varepsilon^{*})\geq V_{\text{garb}}(0)$. ∎
Lemma 10 (More information increases total welfare relative to agent optimal garbling).
$W_{\text{garb}}^{\prime}(\varepsilon)\geq V_{\text{garb}}^{\prime}(\varepsilon)$ for all $\varepsilon\in[0,1]$. The inequality is strict if $\Pi_{\text{garb}}^{\prime}(0)>0$, which is true under strict concavity of $\Pi_{0}(p),\Pi_{1}(p)$ (Assumption 1) and the MLRP (Assumption 3).
Proof.
Lemma 6 implies that $\Pi_{\text{garb}}^{\prime}(\varepsilon)>0$, which implies that $W_{\text{garb}}^{\prime}(\varepsilon)\geq V_{\text{garb}}^{\prime}(\varepsilon)$. The inequality is strict under the conditions of Lemma 7. ∎